### v2

# Modulo arithmetic of function spaces: Subset hyperspaces as quotients of function spaces

Abstract.
Let $X$ be a (topological) space and $Cl(X)$ the collection of nonempty closed subsets of $X$. Given a topology on $Cl(X)$, making $Cl(X)$ a space, a (subset) hyperspace of $X$ is a subspace $\mathcal{J}\subset Cl(X)$ with an embedding $X\hookrightarrow\mathcal{J}$, $x\mapsto\{x\}$. In this note, we characterize certain hyperspaces $\mathcal{J}\subset Cl(X)$ as explicit quotient spaces of function spaces $\mathcal{F}\subset X^{Y}$ and discuss metrization of associated compact-subset hyperspaces in this setting. In particular, we find that any hyperspace topology containing the Vietoris topology is a quotient of a function space topology containing the topology of pointwise convergence.
Key words and phrases:
Subset hyperspace, function space, quotient topology, swrc-topology.
2020 Mathematics Subject Classification:
Primary 54B20 54C35; Secondary 54B15 54E35 54A10 54C50
1. Introduction
The phrase “modulo arithmetic” in the title of this paper is based on the understanding that the operation of taking quotients of (algebraic or geometric) structures by substructures to obtain new structures might be seen as generalizing the modulo arithmetic of integers, with quotient rings being the closest generalizations of the ring $\mathbb{Z}_{n}:=\mathbb{Z}/n\mathbb{Z}$ of modulo $n$ integers. Our situation in particular admits the simplified formula
(subset hyperspace) = (function space) modulo (quotient map).
Theorem 3.3’s Remark (5)-(7) makes a connection with quotients of groups, rings, and modules.
Preliminary notation and terminology
Given a set $X$ and a collection $\mathcal{B}\subset\mathcal{P}(X)$ of subsets of $X$, the topology on $X$ generated by $\mathcal{B}$, i.e., with $\mathcal{B}$ as a subbase, is denoted by $\langle\mathcal{B}\rangle$. Given two (topological) spaces $X_{1}$ and $X_{2}$, $X_{1}\cong X_{2}$ means $X_{1}$ and $X_{2}$ are homeomorphic. Let $X=(X,\tau)$ be a space. A topology $\tau^{\prime}\subset\mathcal{P}(X)$ on $X$ is $\tau$-compatible if $\tau\subset\tau^{\prime}$ (where $\tau^{\prime}$ can therefore inherit certain desirable properties of $\tau$, e.g., if $\tau$ is $T_{0}$, $T_{1}$, or $T_{2}$ respectively, then so is $\tau^{\prime}$). If $A\subset X$, the closure of $A$ in $X$ is denoted by $cl_{X}(A)$, or by $\overline{A}$ if the underlying space $X$ is understood. We say that $A\subset X$ is precompact in $X$ if $cl_{X}(A)$ is compact in $X$. Let $Z$ be a metric space. The completion of $Z$ is (up to isometry) the complete metric space $\widetilde{Z}$ containing $Z$ as a dense subset. We say $A\subset Z$ is totally bounded if $A$ is precompact in $\widetilde{Z}$. It is easy to verify that a set $A\subset Z$ is totally bounded in the usual sense if and only if $A$ is precompact in $\widetilde{Z}$. It might be worth noting that in the context of [akofor2024], “precompact” is automatically equivalent to “totally bounded”, as explained by [akofor2024, Notation 1].
Let $\mathcal{H}(X)$ denote the class of homeomorphisms of $X$ and let $\mathcal{C}\subset\mathcal{H}(X)$. By geometry of $X$ (resp., $\mathcal{C}$-geometry of $X$) we mean the study of one or more properties of $X$ that are invariant under specific homeomorphisms of $X$ (resp., homeomorphisms of $X$ in $\mathcal{C}$), where the invariant properties are accordingly called geometric properties (resp., $\mathcal{C}$-geometric properties) of $X$. By metric geometry we mean geometry that employs metrics. With $\mathcal{P}^{\ast}(X)$ denoting the set of nonempty subsets of $X$, we would like to study the simplest kinds of topologies on $\mathcal{J}\subset\mathcal{P}^{\ast}(X)$ using better understood topologies on $\mathcal{F}\subset X^{Y}$ (for sets $Y$). Since a $T_{1}$ space $X$ can be seen as a subset of $\mathcal{P}^{\ast}(X)$ in a natural way (through the inclusion $X\hookrightarrow\mathcal{P}^{\ast}(X),~x\mapsto\{x\}$), we often consider those topologies on $\mathcal{J}\supset\textrm{Singlt}(X):=\big\{\{x\}:x\in X\big\}$ that can be seen as extensions of the topology of $X$. Such topologies are called hypertopologies of $X$, and the associated spaces $\mathcal{J}$, $\textrm{Singlt}(X)\subset\mathcal{J}\subset\mathcal{P}^{\ast}(X)$, are called (subset) hyperspaces of $X$. A closed-subset hyperspace, compact-subset hyperspace, or bounded-subset hyperspace of $X$ is a hyperspace consisting respectively of closed subsets, compact subsets, or bounded subsets of $X$.
Let $X$ be a $T_{1}$ space, $Cl(X)\subset\mathcal{P}^{\ast}(X)$ the set of nonempty closed subsets of $X$ as a hyperspace whose topology is comparable to the Vietoris topology $\tau_{v}$, $K(X)\subset Cl(X)$ the subspace consisting of all nonempty compact subsets of $X$, and $FS_{n}(X):=\{A\in K(X):1\leq|A|\leq n\}$ the subspace consisting of all nonempty finite subsets of $X$ of cardinality at most $n$. When $X$ is a metric space, we further let $BCl(X)\subset Cl(X)$ denote the subspace consisting of all nonempty bounded closed subsets of $X$. Let $Y$ be a set. On relevant sets of functions $\mathcal{F}\subset X^{Y}$, besides the standard product topology (or topology of pointwise convergence) $\tau_{p}$ and related topologies, we will introduce a preferred topology $\tau_{\pi}$ on $\mathcal{F}\subset X^{Y}$ (Definition 7) in order to discuss metrization of compact-subset hyperspaces (Theorems 4.6 and 4.9). Since a closed set $C\in Cl(X)$ can be seen as the closure of the image $f(Y)$ of some function $f\in X^{Y}$, the image-closure assignment (named unordering map in Definition 3) given by
(1) $q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau_{\pi q}),~f\mapsto cl_{X}(f(Y))$
induces a quotient topology $\tau_{\pi q}$, the $\tau_{\pi}$-quotient topology (footnote [^1] That is, $\tau_{\pi q}:=\sup\big\{\tau~|~q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau)~\textrm{is surjective and continuous}\big\}$ is the largest topology on $Cl_{Y}(X)$ with respect to which the map $q:(\mathcal{F},\tau_{\pi})\rightarrow Cl_{Y}(X)$ is surjective and continuous. ), on the $Y$-indexed closed subsets
(2) $Cl_{Y}(X):=q(\mathcal{F}).$
Motivation
First, we would like to have a straightforward description (by means of a natural quotient map) of the relation between basic function space topologies (e.g., the topologies of pointwise convergence, compact convergence, uniform convergence) and the Vietoris topology of subset hyperspaces. We expect this to facilitate the use of function spaces to study subset hyperspaces and vice versa.
Second, for a metric space $X$, we seek a suitable geometric framework for answering [akofor2024, Question 5.1] concerning characterization/representation of Lipschitz paths in $BCl(X)$ in terms of Lipschitz paths in $X$ or $\widetilde{X}$ (as indicated in Question 5). Due to the key role of Hausdorff distance in the results of [akofor2024], the desired geometric framework appears to require a deeper understanding of the relationship between certain basic function space topologies and the metrization of associated hypertopologies by the Hausdorff distance $d_{H}$ (Equation (3)).
Our work is therefore primarily motivated by the need to use metrizability of certain hyperspaces $\mathcal{J}\subset K(X)$, viewed as quotients of function spaces $\mathcal{F}\subset X^{Y}$ (suitably topologized), to strengthen our understanding/interpretation of $d_{H}$ as a key metric on $\mathcal{J}$. Our main results (Theorems 4.2, 4.6, 4.9, and 5.1) are based on the observation (via Lemma 3.1, Lemma 3.2, and Theorem 3.3) that basic hyperspace topologies (with reasonable separation properties) admit a natural description as quotients of basic function space topologies (with reasonable separation properties).
Summary and highlight of main results
To describe our main results, consider the following question.
  1. Question A: Let $X$ be a space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$ a subset satisfying $q(\mathcal{F})=q(X^{Y})$. Define $Cl_{Y}(X):=q(X^{Y})$ and $K_{Y}(X):=Cl_{Y}(X)\cap K(X)$, where $q$ is the map given by
    $q:\mathcal{F}\rightarrow Cl_{Y}(X),~~f\mapsto cl_{X}(f(Y)).$
  1. (1)
    Can we choose and characterize a topology $\tau_{\pi}\supset\tau_{p}$ on $\mathcal{F}$ such that the following hold?
    1. (a)
      The map $q:(\mathcal{F},\tau_{\pi})\rightarrow Cl_{Y}(X)$ induces a quotient topology $\tau_{\pi q}\supset\tau_{v}$ on $Cl_{Y}(X)$.
    2. (b)
      The compact-subset hyperspace $(K_{Y}(X),\tau_{\pi q})$ is metrizable whenever $X$ is metrizable.
  2. (2)
    If $\tau\supset\tau_{v}$ is a topology on $Cl_{Y}(X)$, does there exist a topology $\widetilde{\tau}\supset\tau_{p}$ on $\mathcal{F}$ such that $\tau\supset\widetilde{\tau}_{q}\supset\tau_{v}$ (where $\widetilde{\tau}_{q}$ is the quotient topology induced on $Cl_{Y}(X)$ by $\widetilde{\tau}$ via the map $q$)?
In Section 3, we describe a class of hyperspaces as quotients of function spaces, noting that some related work has been considered in [bartsch2014] (see Fact 1 below). Next in Section 4, we discuss the concrete realization of certain preferred function space topologies and metrization of compact-subset hyperspaces, and (under some regularity constraints on $\mathcal{F}$) give a positive answer to Question A (1a)&\&(1b) in Theorems 4.2 &\& 4.9 respectively. In Section 5, we discuss the realization of $\tau_{v}$-compatible hyperspace topologies as quotients of $\tau_{p}$-compatible function space topologies, and give a positive answer to Question A(2) in Theorem 5.1.
Fact 1 (Definition 8’s Remark (1)).
For any $T_{3}$-space (i.e., regular Hausdorff space) $X$, according to [bartsch2014, Theorem 2.4], there always exists a compact space $Y$ such that we have a quotient map
$q:C(Y,X)\subset(X^{Y},\tau_{\textrm{co}})\rightarrow(K(X),\tau_{v}),$
where $C(Y,X):=\{\textrm{continuous}~f\in X^{Y}\}$ and $\tau_{co}$ is the compact-open topology (Definition 8).
Fact 1 may be seen as a special case of Theorem 5.1 (on the existence of a $\tau_{p}$-compatible topology on $\mathcal{F}\subset X^{Y}$ with a $\tau_{v}$-compatible quotient). We conclude in Section 6 with some interesting questions.
Throughout, we appeal to intuition by preferably employing sequences and nets (instead of open sets) in our results and proofs whenever this seems convenient.
2. Preliminary remarks
Subset hyperspaces have been studied as function spaces in [Render1993, McCoy1998, DolMyn2010], and as quotient spaces of function spaces in [bartsch2014, bartsch2004, bartsch2002]. These works mainly characterize and compare various geometric properties of those restrictions and extensions of the Vietoris topology that arise through classical (operations like embedding and compactification around) function space topologies. Our discussion is more focused towards explicitly constructing certain generalizations of the pointwise convergence topology $\tau_{p}$ (of function spaces) whose quotients agree with the metrizable Vietoris topology $\tau_{v}$ of compact-subset hyperspaces.
One of our main results, namely, metrization of compact-subset hyperspaces in Theorem 4.9, is well-known for standard hyperspaces with the Vietoris topology (see Lemma 4.5, [IllanNadl1999, Theorems 2.4, 3.1, 3.2, 3.4], [michael1951, Theorem 4.9.13], and [DolMyn2010, Bart2014, McCoy1998, RodRom2002, LeviEtal1993, Render1993]). Our goal here is simply to present a specialized review that emphasizes the description of $\tau_{v}$-compatible subset hyperspaces as quotients of $\tau_{p}$-compatible function spaces (where the role originally played by the Vietoris topology $\tau_{v}$ is now played by $\tau_{v}$-compatible quotient topologies induced by relevant function space topologies). The discussion may be viewed as an extension of preliminary discussions for the case of finite-subset hyperspaces in [akofor2020, Chapter 1, especially around Definition 1.0.1 and Proposition 1.2.2].
Definition 1 (Hausdorff distance).
Let $X$ be a metric space, $x\in X$, $A\subset X$, $\varepsilon>0$,
$\textstyle d(x,A):=\inf\limits_{a\in A}d(x,a),$
$\textstyle N_{\varepsilon}(A):=\{x\in X:d(x,A)<\varepsilon\},~\textrm{and}$~
$\overline{N}_{\varepsilon}(A)=A_{\varepsilon}:=\{x\in X:d(x,A)\leq\varepsilon\}.$
If $A,B\in BCl(X)$, the Hausdorff distance between $A$ and $B$ is
(3) $d_{H}(A,B)$ $\textstyle:=\max\{\sup_{a\in A}d(a,B),\sup_{b\in B}d(b,A)\}$
$=\inf\left\{r>0:A\cup B\subset\overline{N}_{r}(A)\cap\overline{N}_{r}(B)\right\}$
$=\textstyle\sup_{x\in A\cup B}|d(x,A)-d(x,B)|$
$\textstyle=\sup_{x\in X}|d(x,A)-d(x,B)|.$
Definition 2 (Saturated set).
Let $g:A\rightarrow B$ be a map. A subset $S\subset A$ is $g$-saturated if $g^{-1}(g(s))\subset S$ for every $s\in S$ (i.e., $g^{-1}(g(S))=S$, or equivalently, $S=g^{-1}(T)$ for some $T\subset B$).
Recall that a quotient map $f:(U,\tau_{U})\rightarrow(V,\tau_{V})$ makes $V$ a quotient (or quotient space) of $U$, in which case we will call $\tau_{V}$ the $f$-quotient topology induced on $V$ by $\tau_{U}$, written as
$\tau_{V}=\tau_{Uf},$
which is a prelude to the notation introduced in Item 4 of Definition 3.
Remark.
Let $q:U\rightarrow V$ be a quotient map.
  1. (1)
    Given $I\subset U$, the restriction $q|_{I}:I\rightarrow q(I)$ need not be a quotient map with respect to the subspace topologies. Indeed, if $P$ is open in $V$, i.e., $E:=P\cap q(I)$ is open in $q(I)$, then $q|_{I}^{-1}(E):=q^{-1}(E)\cap I=q^{-1}(P)\cap I$ (which is open in $I$), but $q^{-1}(F)\cap I=q^{-1}(P)\cap I$ (for some $F\subset V$) does not always imply that $F\cap q(I)$ is open in $q(I)$.
  2. (2)
    If $O\subset U$ (resp., $C\subset U$) is a $q$-saturated open (resp., closed) set, then $q|_{O}:O\rightarrow q(O)$ (resp., $q|_{C}:C\rightarrow q(C)$) is a quotient map with respect to the subspace topologies. Indeed, in Part (1) above with $I$ replaced by $O$, $q^{-1}(F)\cap O=q^{-1}(P)\cap O$ ($\iff$ $F\cap q(O)=P\cap q(O)$) is open in $U$ if and only if $F\cap q(O)$ is open in $q(O)$.
  3. (3)
    In particular, if $O\subset V$ (resp., $C\subset V$) is an open (resp., closed) set, then $q^{-1}(O)\subset U$ (resp., $q^{-1}(C)\subset U$) is a $q$-saturated open (resp. closed) set, giving a quotient map
    $q|_{q^{-1}(O)}:q^{-1}(O)\rightarrow O,~~(\textrm{resp.,}~~q|_{q^{-1}(C)}:q^{-1}(C)\rightarrow C).$
Due to (1) above, a quotient map $q:(X^{Y},\tau)\rightarrow(q(X^{Y}),\tau^{\prime})$ need not automatically restrict to a quotient map $q:(\mathcal{I},\tau)\subset(X^{Y},\tau)\rightarrow(q(\mathcal{I}),\tau^{\prime})\subset(q(X^{Y}),\tau^{\prime})$. Consequently, we will typically (i) fix a relevant subset $\mathcal{F}\subset X^{Y}$ satisfying $q(\mathcal{F})=q(X^{Y})$ and (ii) directly specify quotient maps
$q:(\mathcal{F},\tau)\subset X^{Y}\rightarrow(q(X^{Y}),\tau_{q}),$
which need not be restrictions of quotient maps $(X^{Y},\tau^{\prime})\rightarrow(q(X^{Y}),\tau^{\prime}_{q})$, even if $(\mathcal{F},\tau)\subset(X^{Y},\tau)$ (i.e., the topology of $\mathcal{F}$ happens to be a subspace topology of a topology on $X^{Y}$). (footnote [^2] If AA is a set and $(B,\tau)$ a space such that $A\subset B$, then the subspace topology on $A$ is the topology $\tau^{i}$ that makes the inclusion $i:(A,\tau^{i})\hookrightarrow(B,\tau)$ a quotient map, where we know $\tau^{i}=\tau\cap A:=\{U\cap A:U\in\tau\}$. Similarly, if $(A,\tau)$ is a space and $B$ a set such that $A\subset B$, then the inclusion $i:A\hookrightarrow B$ induces the $i$-quotient topology (call it the superspace topology) $\tau_{i}$ on $B$ given by $\tau_{i}=\{U\subset B:U\cap A\in\tau\}$. )
3. Subset hyperspaces as quotients of function spaces
We will characterize a class of subset hyperspaces as explicit quotients of function spaces. To have reasonable separation properties, we aim to choose the function space topologies to be $\tau_{p}$-compatible and likewise choose the hyperspaces topologies to be $\tau_{v}$-compatible.
Definition 3 (Indexed subset hyperspaces: Limit Vietoris topology, Unordering map).
Let $Y$ be a set. Given a family of spaces $\mathcal{X}=\{X_{y}:y\in Y\}$, let $X=\bigcup\mathcal{X}:=\bigcup_{y\in Y}X_{y}$, and let
$\prod\mathcal{X}=\prod_{y\in Y}X_{y}:=\{\textrm{maps}~f:Y\rightarrow X,~y\mapsto f_{y}\in X_{y}\}=\{(f_{y})_{y\in Y}:f_{y}\in X_{y}\}$
be their Cartesian product as sets. Recall that the product topology $\tau_{p}$ on $\prod\mathcal{X}$} has base sets
$[O_{F}]_{p}=[\{O_{y}:y\in F\}]_{p}:=\big\{f\in{\prod\mathcal{X}}:f_{y}\in O_{y}~\forall y\in F\big\}=\prod_{y\in F}O_{y}~\times~\prod_{y\in Y\backslash F}X_{y},$
for finite subsets $F\subset Y$ and open subset collections $O_{F}=\{O_{y}\subset X_{y}:y\in F\}$.
[…]
$[T_{F}]_{v}=[\{T_{\alpha}:{\alpha}\in F\}]_{v}$ $:=\textstyle\{A\in\mathcal{P}^{\ast}(X):A\subset\bigcup_{\alpha}T_{\alpha},~A\cap T_{\alpha}\neq\emptyset~\forall{\alpha}\in F\},$
for finite collections $T_{F}=\{T_{\alpha}:{\alpha}\in F\}\subset\mathcal{O}(X)$, where $F$ is an arbitrary finite set. (footnote [^3] The collection $\mathcal{B}:=\{[T_{F}]_{v}:F~\textrm{finite}\}$ indeed forms a base for a topology on $\mathcal{P}^{\ast}(X)$, because if we let $T_{F}=\{T_{1},...,T_{n}\}$ and $T^{\prime}_{F^{\prime}}=\{T^{\prime}_{1},...,T^{\prime}_{n^{\prime}}\}$, then $T_{F}]_{v}\cap[T^{\prime}_{F^{\prime}}]_{v}=\big[\{(\bigcup_{i}T_{i})\cap T_{1}^{\prime},\cdots,(\bigcup_{i}T_{i})\cap T_{n^{\prime}}^{\prime},T_{1}\cap(\bigcup_{j}T^{\prime}_{j}),\cdots,T_{n}\cap(\bigcup_{j}T^{\prime}_{j})\}\big]_{v}=\big[\big\{(\bigcup T_{F})\cap T^{\prime}_{\alpha^{\prime}}:\alpha^{\prime}\in F^{\prime}\big\}\cup\big\{T_{\alpha}\cap\bigcup T^{\prime}_{F^{\prime}}:\alpha\in F\big\}\big]_{v}$, which also lies in $\mathcal{B}$. ).
Let us now define the following items (where items (1)-(4) have already appeared in the introduction under a less general setup):
  1. (1)
    The set of nonempty closed subsets of $X$: $Cl(X)\subset\mathcal{P}^{\ast}(X)$.
  2. (2)
    The set of nonempty compact subsets of $X$: $K(X)\subset Cl(X)$.
  3. (3)
    The set of nonempty bounded closed subsets of $X$ (if $X$ is a metric space): $BCl(X)\subset Cl(X)$.
  4. (4)
    The unordering map $q:{\prod\mathcal{X}}\rightarrow Cl(X),~f\mapsto cl_{X}(f(Y))$.
    Given a $j$-labeled topology $\tau_{j}$ on $\mathcal{F}\subset{\prod\mathcal{X}}$, we denote by $\tau_{jq}$ the topology (called the $\tau_{j}$-quotient topology) on $q(\mathcal{F})\subset Cl(X)$ induced by the restriction $q:(\mathcal{F},\tau_{j})\rightarrow q(\mathcal{F})=(q(\mathcal{F}),\tau_{jq})$ as a quotient map. (footnote [^4] Recall that $\tau_{jq}$ is the strongest topology $\tau_{jc}$ on $q(\mathcal{F})\subset{\prod\mathcal{X}}$ such that $q:(\mathcal{F},\tau_{j})\rightarrow(q(\mathcal{F}),\tau_{jc})$ is continuous. In particular, if $\tau_{j}:=\tau\cap\mathcal{F}$ (i.e., $\tau_{j}$ is the $\tau$-subspace topology of $\mathcal{F}$ associated with some topology $\tau$ on $\prod\mathcal{X}$}) then $\tau_{q}\cap q(\mathcal{F})\subset\tau_{jq}=(\tau\cap\mathcal{F})_{q}$ (where $\tau_{q}$ is the $\tau$-quotient topology of $q({\prod\mathcal{X}})$ and $\tau_{q}\cap q(\mathcal{F})$ is the $\tau_{q}$-subspace topology of $q(\mathcal{F})\subset q({\prod\mathcal{X}})$). Therefore, continuity of $q:(\mathcal{F},\tau\cap\mathcal{F})\rightarrow(q(\mathcal{F}),(\tau\cap\mathcal{F})_{q})$ implies continuity of $q:(\mathcal{F},\tau\cap\mathcal{F})\rightarrow(q(\mathcal{F}),\tau_{q}\cap q(\mathcal{F}))$. )
  5. (5)
    The set of $Y$-indexed closed subsets of $X$: $Cl_{Y}(\mathcal{X}):=Cl(X)\cap q({\prod\mathcal{X}})\stackrel{{\scriptstyle(\ast)}}{{=}}q({\prod\mathcal{X}})$ (where step ($\ast$) is due to the use of closure in the definition of $q$). Also, let
    $Cl(Y,\mathcal{X}):=q^{-1}(Cl_{Y}(\mathcal{X}))\stackrel{{\scriptstyle(\ast)}}{{=}}{\prod\mathcal{X}}.$
  6. (6)
    The set of $Y$-indexed compact subsets of $X$: $K_{Y}(\mathcal{X}):=K(X)\cap q({\prod\mathcal{X}})$. Also, let
    $K(Y,\mathcal{X}):=q^{-1}(K_{Y}(\mathcal{X}))~~\stackrel{{\scriptstyle\textrm{$q$-saturated}}}{{\subset}}~~{\prod\mathcal{X}}.$
  7. (7)
    The set of $Y$-indexed bounded closed subsets of $X$ (if $X$ is a metric space): $BCl_{Y}(\mathcal{X}):=BCl(X)\cap q({\prod\mathcal{X}})$. Also, let
    $BCl(Y,\mathcal{X}):=q^{-1}(BCl_{Y}(\mathcal{X}))~~\stackrel{{\scriptstyle\textrm{$q$-saturated}}}{{\subset}}~~{\prod\mathcal{X}}.$
  8. (8)
    The set of nonempty finite subsets of $X$: $FS(X):=\{A\in K(X):|A|<\infty\}$.
  9. (9)
    The set of nonempty $n$-finite subsets of $X$: $FS_{n}(X):=\{A\in FS(X):|A|\leq n\}$.
When $\mathcal{X}=\{X\}$, i.e., $X_{y}=X$ for all $y\in Y$, we will replace $\mathcal{X}$ with $X$ in the $Y$-indexed subset/function spaces above by setting
$Cl_{Y}(X):=Cl_{Y}(\mathcal{X}),~~Cl(Y,X):=Cl(Y,\mathcal{X}),~~K_{Y}(X):=K_{Y}(\mathcal{X}),~~K(Y,X):=K(Y,\mathcal{X}),$
$BCl_{Y}(X):=BCl_{Y}(\mathcal{X}),~~BCl(Y,X):=BCl(Y,\mathcal{X}).$
Note 1.
If $X$ is $T_{1}$ (i.e., singletons of $X$ are closed in $X$) then $X\cong F_{1}(X)=\textrm{Singlt}(X)$.
Definition 4 ($q$-full subset).
In the setup of Definition 3, a subset $\mathcal{F}\subset\prod\mathcal{X}$ is $q$-full if
$q(\mathcal{F})=q(\prod\mathcal{X}).$
Notice that if $X,Y$ are sets and $\mathcal{F}\subset X^{Y}$ is $q$-full, then we have an injection $X\hookrightarrow\mathcal{F},~x\mapsto c_{x}$, where $c_{x}$ is the constant map $Y\rightarrow X,~y\mapsto x$. Consequently, if $X$ is a space, then topologies on $\mathcal{F}$ (just like hypertopologies of a $T_{1}$ space $X$) may be seen as extensions/generalizations of the topology of $X$.
Definition 5 (rc-topology, rc-space).
In the setup of Definition 3, fix a $q$-full subset $\mathcal{F}=\mathcal{F}(Y,\mathcal{X})\subset\prod\mathcal{X}$ and consider a topology $\tau$ on $\mathcal{F}$. A net $f_{\alpha}\in(\mathcal{F},\tau)$ is compactly-ranged (hence a cr-net) if there exists a compact set $K\subset X=\bigcup_{y\in Y}X_{y}$ and a tail $T_{\beta}:=\{f_{\alpha}:\alpha\geq\beta\}$ of $f_{\alpha}$ such that $\bigcup q(T_{\beta}):=\bigcup_{\alpha\geq\beta}q(f_{\alpha})\subset K$.
The topology $\tau$ on $\mathcal{F}$ is range-compact (hence an rc-topology on $\mathcal{F}$, making $(\mathcal{F},\tau)$ an rc-space) if every cr-net $f_{\alpha}\in(\mathcal{F},\tau)$ has a convergent subnet.
Remark.
Let $X$ be a space and $Y$ a set. Then, by Tychonoff’s product theorem, $(X^{Y},\tau_{p})$ is an rc-space: Indeed, if $K\subset X$ is compact, then $K^{Y}\subset(X^{Y},\tau_{p})$ is a compact subspace. We also recall that in a compact space, a sequence (being a net) has a convergent subnet, but not necessarily a convergent subsequence, i.e., a compact space need not be sequentially compact.
Definition 6 (wrc-topology, wrc-space).
In the setup of Definition 3, let $\tau^{\prime}$ be a topology on $Cl_{Y}(\mathcal{X})$, and fix a $q$-full subset $\mathcal{F}=\mathcal{F}(Y,\mathcal{X})\subset\prod\mathcal{X}$. A topology $\tau$ on $\mathcal{F}$ is $\tau^{\prime}$-weakly range-compact (hence a $\tau^{\prime}$-wrc-topology on $\mathcal{F}$, making $(\mathcal{F},\tau)$ a $\tau^{\prime}$-wrc-space) if for any $\tau^{\prime}$-convergent net $C_{\alpha}\in(K_{Y}(\mathcal{X}),\tau^{\prime})$ such that $\bigcup_{\alpha\geq\beta}C_{\alpha}\subset K$ for a compact set $K\subset X$ and some index $\beta$, every net $g_{\alpha}\in q^{-1}(C_{\alpha})$, $\alpha\geq\beta$, has a $\tau$-convergent subnet in $(\mathcal{F},\tau)$. (footnote [^5] Observe that the rc-property is stronger than (i.e., implies) the wrc-property. That is, every rc-topology (such as $\tau_{p}$) is a wrc-topology. ).
A $\tau_{v}$-wrc-topology (resp., $\tau_{v}$-wrc-space) will simply be called a wrc-topology (resp., wrc-space).
Definition 7 (swrc-topology, swrc-space, Standard topology of the indexed subset hyperspaces).
In the setup of Definition 3, fix a $q$-full subset $\mathcal{F}=\mathcal{F}(Y,\mathcal{X})\subset\prod\mathcal{X}$. A wrc-topology $\tau_{\pi}$ on $\mathcal{F}$ is a standard wrc-topology (hence a swrc-topology on $\mathcal{F}$, making $(\mathcal{F},\tau_{\pi})$ a swrc-space) if $\tau_{v}\subset\tau_{\pi q}$ in $K_{Y}(\mathcal{X})$ (where $\tau_{\pi q}$ denotes the $\tau_{\pi}$-quotient topology on $Cl_{Y}(\mathcal{X})$), i.e., if with
$K\mathcal{F}(Y,\mathcal{X}):=\mathcal{F}\cap K(Y,\mathcal{X})~\subset~\bigcup_{K\in K(X)}K^{Y},$
the map
$q|_{K\mathcal{F}(Y,\mathcal{X})}:K\mathcal{F}(Y,\mathcal{X})\subset(\mathcal{F},\tau_{\pi})\rightarrow\big(Cl_{Y}(\mathcal{X}),\tau_{v})$
is continuous.
We will give $Cl_{Y}(\mathcal{X})$ the $\tau_{\pi}$-quotient topology $\tau_{\pi q}$ (as our standard topology on $Cl_{Y}(\mathcal{X})$).
Question 1.
Let $X$ be a space, $Y$ a set (and let $\mathcal{X}:=\{X\}$), and $\mathcal{F}\subset X^{Y}$ a $q$-full subset. How do we explicitly specify the preferred topology $\tau_{\pi}$ of Definition 7? Theorem 4.2 specifies $\tau_{\pi}$ for some special cases where $\mathcal{F}$ is sufficiently well-behaved.
Definition 8 (Set-open topology, Compact-open topology).
Let $X$ be a space, $Y$ a set, and $\mathcal{S}\subset\mathcal{P}(Y)$ a family of subsets of $Y$. The $\mathcal{S}$-open topology $\tau_{\mathcal{S}}$ on $X^{Y}$ is the topology with a subbase given by the sets
$S,O]_{\mathcal{S}}:=\{f\in X^{Y}:f(S)\subset O\},~~~~\textrm{for sets $S\in\mathcal{S}$ and open sets $O\subset X$}.$
In particular, if $Y$ is a space and $\mathcal{S}=K(Y)$, then $\tau_{\mathcal{S}}$ is called the compact-open topology, $\tau_{co}$, on $X^{Y}$.
Remark.
  1. (1)
    For any $T_{3}$-space (i.e., regular Hausdorff space) $X$, according to [bartsch2014, Theorem 2.4], there always exists a compact space $Y$ such that $q\big(C(Y,X)\big)=K(X)$ (where $C(Y,X):=\{\textrm{continuous}~f\in X^{Y}\}$) and we have a quotient map
    $q:C(Y,X)\subset(X^{Y},\tau_{\textrm{co}})\rightarrow(K(X),\tau_{v}).$
  2. (2)
    Based on the setup in Definition 7, if $X,Y$ are spaces and $\tau_{\pi}$ is a swrc-topology on a $q$-full $\mathcal{F}\subset X^{Y}$, then we have the quotient map
    $q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau_{\pi q}).$
    In this special case (where both $X$ and $Y$ are spaces), candidates for $\tau_{\pi}$ in Definition 7 are $q^{-1}(\tau_{v})$ and $\tau_{co}$, and it might therefore be reasonable/sufficient (regarding Question 1) to search for a $\tau_{p}$-compatible $\tau_{\pi}$ in the range (footnote [^6] Consider the map $q:(X^{Y},\tau_{1},\tau_{2})\rightarrow(Cl_{Y}(X),\tau_{1q},\tau_{2q})$. If $\tau_{1}\subset\tau_{2}$, then $B\in\tau_{1q}$ $\iff$ $q^{-1}(B)\in\tau_{1}\subset\tau_{2}$, $\Rightarrow$ $q^{-1}(B)\in\tau_{2}$, $\iff$ $B\in\tau_{2q}$ (i.e., $\tau_{1}\subset\tau_{2}$ $\Rightarrow$ $\tau_{1q}\subset\tau_{2q}$). )
    $q^{-1}(\tau_{v})\subset\langle\tau_{p}\cup\tau_{s}\rangle\subset\tau_{\pi}\subset\tau_{co}$
    (in accordance with   $\tau_{v}\subset\langle\tau_{p}\cup\tau_{s}\rangle_{q}\subset\tau_{\pi q}\subset\tau_{co\!\!\!~q}$   due to Theorem 3.3 and [bartsch2014, Corollary 2.3]), where $\tau_{s}$ is the topology of symmetric convergence (Definition 10).
    In addition to Theorem 4.2, ways of choosing $\tau_{\pi}$ might be found in [bartsch2014, Lemma 2.2], [bartsch2004, Corollary 3.18], and [bartsch2002], and potentially involve generalizations of Tychonoff’s product theorem (e.g., in [park1975, FoxMor1974, kimber1974]).
We will now discuss the Vietoris topology (a more general version of which we have already introduced in Definition 3) in sufficient detail for our subsequent discussion. Using convergence of nets, a more intuitive interpretation of the Vietoris topology is given later in Definition 13.
Definition 9 (Vietoris topology).
Let $X$ be a space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$ a $q$-full subset. For any open set $O\subset X$ and any finite collection of open subsets $O_{F}=\{O_{\alpha}:{\alpha}\in F\}$ of $X$, let
$[O_{F}]_{v}$ $\textstyle:=\{C\in Cl_{Y}(X):C\subset\bigcup_{{\alpha}\in F}O_{\alpha},~C\cap O_{\alpha}\neq\emptyset,\forall{\alpha}\in F\}$
$\textstyle=\left\{q(f):f\in\mathcal{F},~q(f)\subset\bigcup_{{\alpha}\in F}O_{\alpha},~q(f)\cap O_{\alpha}\neq\emptyset,\forall{\alpha}\in F\right\},$
$O^{+}$ $\textstyle:=[\{O\}]_{v}=\{C\in Cl_{Y}(X):C\subset O\}=\left\{q(f):f\in\mathcal{F},~q(f)\subset O\right\},$
$O^{-}$ $\textstyle:=[\{X,O\}]_{v}=[\{O^{c}\}]_{v}{}^{c}=\{C\in Cl_{Y}(X):C\cap O\neq\emptyset\}=\left\{q(f):f\in\mathcal{F},~q(f)\cap O\neq\emptyset\right\}$
$\textstyle=\left\{q(f):f\in\mathcal{F},~f(Y)\cap O\neq\emptyset\right\}.$
The Vietoris topology $\tau_{v}$ (also see Definition 3) of $Cl_{Y}(X)$ is the topology with base
$\mathcal{B}_{v}=\big\{[O_{F}]_{v}:\textrm{$F$ a finite set, $O_{\alpha}\subset X$ open, $\forall{\alpha}\in F$}\big\}.$
Let the upper Vietoris topology $\tau_{v}^{+}$ (resp., lower Vietoris topology $\tau_{v}^{-}$) be generated by
$S\mathcal{B}_{v}^{+}=\big\{O^{+}:\textrm{$O\subset X$ open}\big\}~~\big(\textrm{resp.},~~S\mathcal{B}_{v}^{-}=\big\{O^{-}:\textrm{$O\subset X$ open}\big\}\big).$
The Vietoris topology then satisfies $\tau_{v}=\langle S\mathcal{B}_{v}^{-}\cup S\mathcal{B}_{v}^{+}\rangle=\langle\tau_{v}^{-}\cup\tau_{v}^{+}\rangle$.
Definition 10 (Topology of symmetric convergence).
Let $X$ be a space and $Y$ a set. The topology of symmetric convergence, $\tau_{s}$, on $X^{Y}$ is the topology with base sets
$O]_{s}:=\{f\in X^{Y}:q(f)\subset O\},~~~~\textrm{for open sets}~O\subset X.$
Definition 11 (Finitely $q$-stable set of functions).
Let $X$ be a space and $Y$ a set. A set of functions $\mathcal{F}\subset X^{Y}$ is finitely $q$-stable if for any $f\in\mathcal{F}$, any finite set $F\subset Y$, and any injection $\sigma:F\rightarrow Y$, there exists $g=g_{F,\sigma}\in\mathcal{F}$ such that $f|_{F}=g|_{\sigma(F)}$ and $q(f)=q(g)$.
Lemma 3.1 (Lower quotient topology).
Let $X$ be a space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$ a q-full finitely $q$-stable subset. Then with respect to $\mathcal{F}$, we have
  1. (i)
    $\tau_{pq}=\tau_{v}^{-}$ in $Cl_{Y}(X)$ (i.e., the map $q:(\mathcal{F},\tau_{p})\rightarrow(Cl_{Y}(X),\tau_{v}^{-})$ is a quotient map) and
  2. (ii)
    $q^{-1}(\tau_{v}^{-})\subsetneq\tau_{p}$ unless $Y$ is a singleton.
Proof.
(i) Proving $\tau_{pq}\subset\tau_{v}^{-}$: A set $\mathcal{A}\subset Cl_{Y}(X)$ is $\tau_{pq}$-open iff $q^{-1}(\mathcal{A})=\left\{f\in\mathcal{F}:q(f)\in\mathcal{A}\right\}$ is $\tau_{p}$-open, i.e., iff there exist a collection of finite sets $\{F_{i}\subset Y\}_{i\in I}$ and open sets $\{O_{y}^{i}\subset X:i\in I,y\in F_{i}\}=\bigcup_{i\in I}\{O_{y}^{i}\subset X:y\in F_{i}\}=\bigcup_{i\in I}O^{i}_{F_{i}}$ such that
$\textstyle q^{-1}(\mathcal{A})$ $\textstyle=\bigcup_{i\in I}[O_{F_{i}}^{i}]_{p}=\bigcup_{i\in I}\{f\in\mathcal{F}:~f(y)\in O_{y}^{i},\forall y\in F_{i}\},$
which is the general form of a $\tau_{p}$-open set in $\mathcal{F}$. By applying $q$ (and noting $q:\mathcal{F}\rightarrow Cl_{Y}(X)$ is surjective, hence $q(q^{-1}(\mathcal{A}))=\mathcal{A}$), we get
$\mathcal{A}$ $\textstyle=q(q^{-1}(\mathcal{A}))=\bigcup_{i\in I}\left\{q(f):f\in\mathcal{F},~f(y)\in O^{i}_{y},\forall y\in F_{i}\right\}$
$\textstyle\stackrel{{\scriptstyle(\ast)}}{{=}}\bigcup_{i\in I}\left\{q(f):f\in\mathcal{F},~f(Y)\cap O_{y}^{i}\neq\emptyset,\forall y\in F_{i}\right\}$
$\textstyle=\bigcup_{i\in I}\left\{q(f):f\in\mathcal{F},~q(f)\cap O_{y}^{i}\neq\emptyset,\forall y\in F_{i}\right\}=\bigcup_{i\in I}\bigcap_{y\in F_{i}}(O_{y}^{i})^{-}$
$\textstyle\in\tau_{v}^{-},$
where at step ($\ast$), $\subset$ is obvious and $\supset$ follows from finite $q$-stability of $\mathcal{F}$.
Proving $\tau_{pq}\supset\tau_{v}^{-}$ (i.e., $q:(\mathcal{F},\tau_{p})\rightarrow(Cl_{Y}(X),\tau_{v}^{-})$ is continuous): Let $O\subset X$ be open. Then
$q^{-1}(O^{-})$ $\textstyle=q^{-1}\left\{q(f):f\in\mathcal{F},q(f)\cap O\neq\emptyset\right\}=\bigcup_{f\in\mathcal{F}}\left\{q^{-1}(q(f)):q(f)\cap O\neq\emptyset\right\}$
$\textstyle=\bigcup_{f\in\mathcal{F}}\left\{g\in\mathcal{F}:q(g)=q(f),q(f)\cap O\neq\emptyset\right\}=\left\{g\in\mathcal{F}:q(g)\cap O\neq\emptyset\right\}$
$\textstyle=\left\{g\in\mathcal{F}:g(Y)\cap O\neq\emptyset\right\}=\left\{g\in\mathcal{F}:\exists y\in Y,g(y)\in O\right\}=\bigcup_{y\in Y}\left\{g\in\mathcal{F}:g(y)\in O\right\}$
$\textstyle=\bigcup_{y\in Y}[(y,O)]_{p}\in\tau_{p}.$
(ii) This follows from the observation that the subbase elements $\{[y,O]_{p}:y\in Y\}$ of $\tau_{p}$ can distinguish the points of $Y$, meanwhile, by construction, neither the elements of $q^{-1}(\tau_{v}^{-})$ nor those of $q^{-1}(\tau_{v}^{+})$ can distinguish the points of $Y$. Indeed, if $|Y|\geq 2$, pick $y_{1},y_{2}\in Y$ such that $y_{1}\neq y_{2}$ and $y_{1},O]_{p}\neq[y_{2},O]_{p}$. Then each $y_{i},O]_{p}$ (an element of $\tau_{p}$) depends asymmetrically on $y_{1}$ and $y_{2}$. But every member of $q^{-1}(\tau^{-}_{v})$ depends symmetrically on $y_{1}$ and $y_{2}$ (as the expression for $q^{-1}(O^{-})$ above shows), and so $q^{-1}(\tau^{-}_{v})$ does not contain $y_{i},O]_{p}$. That is, if $|Y|\geq 2$, then $\tau_{p}\backslash q^{-1}(\tau_{v}^{-})\neq\emptyset$ (provided $X$ is a sufficiently nontrivial space). ∎
Remark.
The following are further observations following the proof of Lemma 3.1.
  1. (1)
    For $y\in Y$ and any open set $O\subset X$, we have both $q([O]_{s})=O^{+}$ and
    $q([(y,O)]_{p})=\left\{q(f):f\in\mathcal{F},f(y)\in O\right\}\stackrel{{\scriptstyle\textrm{finite $q$-stability}}}{{=}}\left\{q(f):f\in\mathcal{F},q(f))\cap O\neq\emptyset\right\}=O^{-}.$
    However, these relations alone do not guarantee openness of $q:(\mathcal{F},\tau)\rightarrow(Cl_{Y}(X),\tau_{v})$ whether for $\tau=\tau_{p}$, $\tau=\tau_{s}$, or $\tau=\langle\tau_{p}\cup\tau_{s}\rangle$, since $q$ (like other maps in general) need not preserve finite intersections.
  2. (2)
    $q^{-1}(\tau_{v}^{+})=\tau_{s}$ and, if $X$ is $T_{1}$, then $\tau_{s}\subset\tau_{p}$ $\iff$ $Y$ is finite, since
    $q^{-1}(O^{+})$ $\textstyle=q^{-1}\left\{q(f):f\in\mathcal{F},q(f)\subset O\right\}=\bigcup\limits_{f\in\mathcal{F}}\left\{q^{-1}(q(f)):q(f)\subset O\right\}=\bigcup\limits_{f\in\mathcal{F}}\left\{g\in\mathcal{F}:q(g)=q(f)\subset O\right\}$
    $\textstyle=\left\{g\in\mathcal{F}:q(g)\subset O\right\}=[O]_{s}=\left\{g\in\mathcal{F}:g(Y)\subset O\right\}\cap\left\{g\in\mathcal{F}:\partial g(Y)\subset O\right\}$
    $\textstyle=\bigcap_{y\in Y}\left\{g\in\mathcal{F}:g(y)\in O\right\}\cap\left\{g\in\mathcal{F}:\partial g(Y)\subset O\right\}=\left[\bigcap_{y\in Y}[(y,O)]_{p}\right]\cap\left\{g\in\mathcal{F}:\partial g(Y)\subset O\right\}$
    $\textstyle\in\tau_{p}~~\iff~~\textrm{$Y$ is finite}.$
  3. (3)
    So, if $X$ is $T_{1}$, then on $\mathcal{F}\subset X^{Y}$ ($X$ nontrivial), $q^{-1}(\tau_{v})\subset\tau_{p} \iff$ $Y$ is finite.
Lemma 3.2 (Upper quotient topology).
Let $X$ be a space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$ a $q$-full subset. We have:
  1. (i)
    $\tau_{sq}=\tau_{v}^{+}$ in $Cl_{Y}(X)$ (i.e., the map $q:(\mathcal{F},\tau_{s})\rightarrow(Cl_{Y}(X),\tau_{v}^{+})$ is a quotient map),
  2. (ii)
    $\tau_{s}=q^{-1}(\tau_{v}^{+})$, and
  3. (iii)
    If $X$ is $T_{1}$, then $\tau_{s}\subset\tau_{p}$ $\iff$ $Y$ is finite (by Lemma 3.1’s Remark (2)).
Proof.
(i) Proving $\tau_{sq}\subset\tau_{v}^{+}$: A set $\mathcal{A}\subset Cl_{Y}(X)$ is $\tau_{sq}$-open iff $q^{-1}(\mathcal{A})=\left\{f\in\mathcal{F}:q(f)\in\mathcal{A}\right\}$ is $\tau_{s}$-open, i.e., iff there exist a collection of open sets $\{O_{i}\subset X:i\in I\}$ such that
$\textstyle q^{-1}(\mathcal{A})$ $\textstyle=\bigcup_{i\in I}[O_{i}]_{s}=\bigcup_{i\in I}\{f\in\mathcal{F}:~q(f)\in O_{i}\},$
which is the general form of a $\tau_{s}$-open set in $\mathcal{F}\subset X^{Y}$. By applying $q$ (and noting $q:\mathcal{F}\rightarrow Cl_{Y}(X)$ is surjective, hence $q(q^{-1}(\mathcal{A}))=\mathcal{A}$) we get
$\mathcal{A}$ $\textstyle=q(q^{-1}(\mathcal{A}))=\bigcup_{i\in I}\left\{q(f):f\in\mathcal{F},~q(f)\subset O_{i}\right\}=\bigcup_{i\in I}O_{i}^{+}\in\tau_{v}^{+}.$
Proving $\tau_{sq}\supset\tau_{v}^{+}$ (i.e., $q:(\mathcal{F},\tau_{s})\rightarrow(Cl_{Y}(X),\tau_{v}^{+})$ is continuous): As in Lemma 3.1’s Remark (2),
$q^{-1}(O^{+})$ $\textstyle=q^{-1}\left\{q(f):f\in\mathcal{F},q(f)\subset O\right\}=\bigcup_{f\in\mathcal{F}}\left\{q^{-1}(q(f)):q(f)\subset O\right\}$
$\textstyle=\bigcup_{f\in\mathcal{F}}\left\{g\in\mathcal{F}:q(g)=q(f)\subset O\right\}$
$\textstyle=\left\{g\in\mathcal{F}:q(g)\subset O\right\}=[O]_{s}\in\tau_{s}.$
(ii) From the above equality, we see that $\tau_{s}=q^{-1}(\tau_{v}^{+})$. ∎
For a finite $Y$, the following theorem (Theorem 3.3) realizes $\tau_{v}$ as a quotient of a $\tau_{p}$-compatible topology. For a general $Y$, the realization of $\tau_{v}$ as a of a $\tau_{p}$-compatible topology will be accomplished in Theorem 5.1 (which requires Theorem 3.3).
Theorem 3.3 ($\tau_{v}$-compatibility of $\langle\tau_{p}\cup\tau_{s}\rangle_{q}$).
Let $X$ be a space, $Y$ a set, $\mathcal{F}\subset X^{Y}$ a $q$-full finitely $q$-stable subset, and consider the map $q:\mathcal{F}\rightarrow Cl_{Y}(X)$. The following are true:
  1. (1)
    $\tau_{v}=\langle\tau_{pq}\cup\tau_{sq}\rangle\subset\langle\tau_{p}\cup\tau_{s}\rangle_{q}$,  where if $X$ is $T_{1}$, then equality holds iff $Y$ is finite (in which case $\tau_{v}=\tau_{pq}$).
  2. (2)
    $q^{-1}(\tau_{v})=\langle q^{-1}(\tau_{v}^{-})\cup q^{-1}(\tau_{v}^{+})\rangle\subset\langle\tau_{p}\cup\tau_{s}\rangle$,  where equality holds iff $Y$ is a singleton.
In particular, the map  $q:(\mathcal{F},\langle\tau_{p}\cup\tau_{s}\rangle)\rightarrow(Cl_{Y}(X),\tau_{v})$  is continuous.
Proof.
By Lemmas 3.1 and 3.2, we get both (1) via $\tau_{v}=\langle\tau_{v}^{-}\cup\tau_{v}^{+}\rangle$ and (2) via $q^{-1}(\tau_{v})=q^{-1}(\langle\tau_{v}^{-}\cup\tau_{v}^{+}\rangle)=\langle q^{-1}(\tau_{v}^{-})\cup q^{-1}(\tau_{v}^{+})\rangle$. ∎
Remark.
Let $X$ be a space, $Y$ a set, and $(\mathcal{F},\tau)\subset X^{Y}$ a $q$-full function space.
  1. (1)
    In the above results for $q:(\mathcal{F},\tau)\rightarrow(Cl_{Y}(X),\tau_{q})$ with $\tau_{q}=\tau_{v}$, if $FS_{n}(X)\subset(Cl_{Y}(X),\tau_{q})$ is closed, then (by Definition 2’s Remark (3)) $q:(q^{-1}(FS_{n}(X))\cap\mathcal{F},\tau)\subset(\mathcal{F},\tau)\rightarrow(FS_{n}(X),\tau_{q})$ is a quotient map as well. In this case, Theorem 3.3(1) holds (i) with $\mathcal{F}$ and $Cl_{Y}(X)$ replaced by $q^{-1}(FS_{n}(X))\cap\mathcal{F}$ and $FS_{n}(X)$ respectively, and (ii) with equality for any $Y$.
  2. (2)
    If $(\mathcal{F},\tau)$ is compact, connected, or path-connected, then so is $(Cl_{Y}(X),\tau_{q})$, since $q:(\mathcal{F},\tau)\rightarrow(Cl_{Y}(X),\tau_{q})$ is continuous.
  3. (3)
    Continuous maps $h:(Cl_{Y}(X),\tau_{q})\rightarrow Z$ (for a space $Z$) are precisely continuous maps $h:(\mathcal{F},\tau)\rightarrow Z$ that are constant on the equivalence classes $f]:=q^{-1}(q(f))=\left\{g\in\mathcal{F}:q(g)=q(f)\right\}$ (i.e., $h|_{[f]}=const$) for all $f\in\mathcal{F}$.
  4. (4)
    Given topologies $\tau_{1}\subset\tau_{2}$ (e.g., topologies of metrics $d_{1}\leq d_{2}$) on $X$, convergence of a net $x_{\alpha}\in(X,\tau_{2})$ (resp., compactness of a set $A\subset(X,\tau_{2})$) implies convergence of $x_{\alpha}\in(X,\tau_{1})$ (resp., compactness of $A\subset(X,\tau_{1})$). That is, the two sets of convergent nets satisfy
    $\tau_{1}\subset\tau_{2}~~\iff~~\textrm{ConvNet}(\tau_{2})\subset\textrm{ConvNet}(\tau_{1}).$
    Similarly, continuity of a map $f\in(X,\tau_{2})^{(Y,\tau_{Y})}$ implies continuity of $f\in(X,\tau_{1})^{(Y,\tau_{Y})}$, i.e.,
    $\tau_{1}\subset\tau_{2}~~\iff~~C\big(Y,(X,\tau_{2})\big)\subset C\big(Y,(X,\tau_{1})\big).$
  5. (5)
    Let $G$ be a topological group (i.e., a group that is a topological space with continuous group-multiplication and inversion) and $H\subset G$ a closed normal subgroup. For each $g\in G$, let $L_{g}:G\rightarrow G,~x\mapsto gx$ be left translation by $g$, and $Lt(H,G):=\{L_{g}|_{H}:g\in G\}\subset G^{H}$. In the anticipated quotient map  $q:Lt(H,G)\rightarrow q(Lt(H,G))\subset Cl_{H}(G),~f\mapsto\overline{f(H)}$,  we have
    $q(Lt(H,G))=\left\{\overline{L_{g}(H)}=gH:g\in G\right\}=G/H.$
  6. (6)
    Let $R$ be a topological ring (i.e., a ring that is a topological space with continuous multiplication and addition) and $I\subset R$ a closed ideal. For each $r\in R$, let $L_{r}:R\rightarrow R,~x\mapsto r+x$ be translation by $r$, and $Lt(I,R):=\{L_{r}|_{I}:r\in R\}\subset R^{I}$. In the anticipated quotient map  $q:Lt(I,R)\rightarrow q(Lt(I,R))\subset Cl_{I}(R),~f\mapsto\overline{f(I)}$,  we have
    $q(Lt(I,R))=\left\{\overline{L_{r}(I)}=r+I:r\in R\right\}=R/I.$
  7. (7)
    Let $R$ be a topological ring and $_{R}M$ a topological $R$-module (i.e., an $R$-module that is a topological space with continuous addition and scalar multiplication) and $N\subset M$ a closed $R$-submodule. For each $m\in M$, let $L_{m}:M\rightarrow M,~x\mapsto m+x$ be translation by $m$, and $Lt(N,M):=\{L_{m}|_{N}:m\in M\}\subset M^{N}$. In the anticipated quotient map  $q:Lt(N,M)\rightarrow q(Lt(N,M))\subset Cl_{N}(M),~f\mapsto\overline{f(N)}$ we have
    $q(Lt(N,M))=\left\{\overline{L_{m}(N)}=m+N:m\in M\right\}=M/N.$
  8. (8)
    Let $Y$ be a manifold, $X=\bigsqcup_{y\in Y}X_{y}=\bigsqcup_{y\in Y}\pi^{-1}(y)$ a fiber bundle over $Y$ with projection $\pi:X\rightarrow Y$. Let $\mathcal{X}:=\{X_{y}\}_{y\in Y}=\{\pi^{-1}(y)\}_{y\in Y}$. Then, in the notation of Definition 3, we get a map on global sections of $X$ given by
    $q:\prod\mathcal{X}\rightarrow Cl_{Y}(\mathcal{X}),~s\mapsto cl_{X}(s(Y)).$
    In this case, the limit topology of Definition 3 on $X$ plays a nontrivial role as a topology (induced by that of the fibers $X_{y}$ of the bundle $X$) that can be compared with the underlying topology of $X$.
4. Concrete quotient-realization and metrization of compact-subset hyperspaces
Concrete quotient-realization of hyperspace topologies
Given a space $X$, a set $Y$, and a $q$-full subset $\mathcal{F}\subset X^{Y}$, we see that $q:(\mathcal{F},q^{-1}(\tau_{v}))\rightarrow(Cl_{Y}(X),\tau_{v})$ is a surjective open continuous map, hence an open quotient map. In general, $q^{-1}(\tau_{v})$ is a highly non-Hausdorff swrc-topology (see Lemma 4.7) and therefore not always a convenient swrc-topology, but can be refined/extended to a $\tau_{p}$-compatible swrc-topology (as shown in Theorem 4.2) if $\mathcal{F}$ meets certain conditions. Meanwhile, the existence of a $\tau_{p}$-compatible (but not necessarily swrc-) topology on $\mathcal{F}$ with a $\tau_{v}$-compatible quotient is proved in Theorem 5.1.
Definition 12 (Chained space, Topologies of uniform/pointwise/compact-uniform convergence).
A chained-space $X=(X,\mathcal{U})$ is a space $X$ together with a chain $\mathcal{U}=(\{\mathcal{U}_{\lambda}\}_{\lambda\in\Lambda},\preceq)$ of open covers of $X$, where $\mathcal{U}_{\lambda}\preceq\mathcal{U}_{\lambda^{\prime}}$ iff for any $O\in\mathcal{U}_{\lambda}$ there is $O^{\prime}\in\mathcal{U}_{\lambda^{\prime}}$ such that $O\subset O^{\prime}$ and [for any $O^{\prime}\in\mathcal{U}_{\lambda^{\prime}}$ and any $\in\mathcal{U}_{\lambda}, OO^{\prime}$ is not a proper subset of OO]”. A collection of open sets $\mathcal{O}$ of $X$ is homogeneous (with respect to $\mathcal{U}$) if $\mathcal{O}\subset\mathcal{U}_{\lambda}$ for some $\lambda$.
Let $X=(X,\mathcal{U})$ be a chained-space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$. A net $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$ converges uniformly on $Z\subset Y$ to $f\in\mathcal{F}$, written $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$, if for any homogeneous system $\mathcal{O}:=\{O_{z}\ni f(z):O_{z}\in\mathcal{U}_{\lambda},z\in Z\}$ of open sets of $X$ ($\big$(e.g., $\mathcal{O}=\{O\}$ for an open neighborhood $\mathcal{U}_{\lambda}\ni O\supset\overline{f(Z)}$ of $\overline{f(Z)}$ in $X$, or equivalently, $q^{-1}(O^{+})\ni f$ of $f$ in $\mathcal{F}\big)$, there exists $\alpha^{\mathcal{O}}\in\mathcal{A}$ such that for each $z\in Z$, $\{f_{\alpha}(z)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O_{z}$ ($\big$(resp., e.g., $\exists\alpha^{O}\in\mathcal{A}$ such that $\bigcup_{\alpha\geq\alpha^{O}}f_{\alpha}(Z)\subset O\big)$. (footnote [^7] We can refer to uniform convergence defined without reference to a chain $\mathcal{U}$ as unconditional-uniform convergence. That is, a net $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$ converges unconditionally-uniformly on $Z\subset Y$ to $f\in\mathcal{F}$, written $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$, if for any system $\mathcal{O}:=\{O_{z}\ni f(z):z\in Z\}$ of open sets in $X$, there exists $\alpha^{\mathcal{O}}\in\mathcal{A}$ such that for each $z\in Z$, $\{f_{\alpha}(z)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O_{z}$.
NB: All subsequent concepts/results based on (conditional-) uniform convergence, that do not refer to (hence do not depend on) any details about the chain $\mathcal{U}$, also apply-to/hold-for unconditional-uniform convergence. In particular, Notes 2-4 below (and more) do not depend on $\mathcal{U}$.
). If a net $\{f_{\alpha}\}\subset\mathcal{F}$ is uniformly convergent on $Y$, then we simply say $\{f_{\alpha}\}$ converges uniformly (or that $\{f_{\alpha}\}$ is a uniformly convergent net), written $f_{\alpha}\stackrel{{\scriptstyle u}}{{\longrightarrow}}f$.
Note 2.
If $Z\subset Z^{\prime}\subset Y$, then $f_{\alpha}\stackrel{{\scriptstyle u|_{Z^{\prime}}}}{{\longrightarrow}}f$ implies $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$ (which holds because each system of open sets $\mathcal{O}:=\{O_{z}\ni f(z):O_{z}\in\mathcal{U}_{\lambda},z\in Z\}$ can be extended to a system of open sets $\mathcal{O}^{\prime}:=\{O_{z}\ni f(z):O_{z}\in\mathcal{U}_{\lambda},z\in Z^{\prime}\}$). In particular, for any $Z\subset Y$,
$\textrm{$f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f~$ ~implies~ $~f_{\alpha}|_{Z}\stackrel{{\scriptstyle\tau_{p}}}{{\longrightarrow}}f|_{Z}~$ (i.e., $~f_{\alpha}(z)\rightarrow f(z)~$ for all $z\in Z$)}.$
A net $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$ converges uniformly with respect to a family of sets $\mathcal{Z}\subset\mathcal{P}^{\ast}(Y)$ to $f\in\mathcal{F}$ (making it a $\mathcal{Z}$-uniformly convergent net), written $f_{\alpha}\stackrel{{\scriptstyle u|_{\mathcal{Z}}}}{{\longrightarrow}}f$, if $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$ for all $Z\in\mathcal{Z}$. We note that $\{f_{\alpha}\}$ is uniformly convergent if and only if $\{Y\}$-uniformly convergent, if and only if (by Note 2 above) uniformly convergent on $Z$ for every $Z\subset Y$.
A subset $\mathcal{C}\subset X^{Y}$ is $\mathcal{Z}$-uniformly closed if every $\mathcal{Z}$-uniformly convergent net in $\mathcal{C}$ converges to a point in $\mathcal{C}$. The topology of $\mathcal{Z}$-uniform convergence $\tau_{uc|_{\mathcal{Z}}}$ on $X^{Y}$ is the topology whose closed sets are the $\mathcal{Z}$-uniformly closed subsets of $X^{Y}$. When $\mathcal{Z}=\{Y\}$, we simply drop $\mathcal{Z}$ from the terminology, i.e., $\mathcal{Z}$-uniformly closed”, $\tau_{uc|_{\mathcal{Z}}}$ (“topology of $\mathcal{Z}$-uniform convergence”), etc become “uniformly closed”, $\tau_{uc}$ (“topology of uniform convergence”), etc. If $\mathcal{Z}=\big\{\{y\}:y\in Y\big\}$, then $\tau_{uc|_{\mathcal{Z}}}=\tau_{p}$, in which case we replace “uniform” or “uniformly” with “pointwise”. When $Y$ is a space and $\mathcal{Z}=K(Y)$, we call $\tau_{\textrm{cc}}\equiv\tau_{\textrm{cuc}}:=\tau_{uc|_{K(Y)}}$ the topology of compact-uniform convergence (which is a generalization of the usual notion of “compact convergence” or “uniform convergence on compact sets” for a metric $X$).
Note 3.
If $\mathcal{Z}\subset\mathcal{Z}^{\prime}$, then $\tau_{uc|_{\mathcal{Z}}}\subset\tau_{uc|_{\mathcal{Z}^{\prime}}}$. To see this, take a $\tau_{uc|_{\mathcal{Z}}}$-closed set $\mathcal{C}\subset X^{Y}$ and show that it is also $\tau_{uc|_{\mathcal{Z}^{\prime}}}$-closed: Indeed, if $f_{\alpha}\in\mathcal{C}$, then $f_{\alpha}\stackrel{{\scriptstyle\tau_{uc|_{\mathcal{Z}^{\prime}}}}}{{\longrightarrow}}f$ implies $f_{\alpha}\stackrel{{\scriptstyle\tau_{uc|_{\mathcal{Z}}}}}{{\longrightarrow}}f$, and so $f\in\mathcal{C}$.
Note 4.
For any $\mathcal{Z}\subset\mathcal{P}(Y)$, with $Y_{\mathcal{Z}}:=\bigcup\{\mathcal{P}(Z):Z\in\mathcal{Z}\}$, we have $\tau_{uc|_{\mathcal{Z}}}=\tau_{uc|_{\{Y_{\mathcal{Z}}\}}}$. This follows from Note 2 above.
Note 5.
Let $X$ be a metric space, $Y$ a space, and $\mathcal{F}\subset X^{Y}$. In $\mathcal{F}$, if $f_{\alpha}\stackrel{{\scriptstyle\tau_{\textrm{cc}}}}{{\longrightarrow}}f$, then $f_{\alpha}\stackrel{{\scriptstyle d^{K}}}{{\longrightarrow}}f$ (i.e., $\forall\varepsilon>0$, $\exists\alpha^{\varepsilon}$ s.t. $d^{K}(f_{\alpha},f)<\varepsilon$ $\forall\alpha\geq\alpha^{\varepsilon}$) for all $K\in K(Y)$, where $d^{K}(f_{\alpha},f):=\sup_{y\in K}d(f_{\alpha}(y),f(y))$. In particular, in the continuous maps $\mathcal{F}:=C(Y,X)\subset X^{Y}$, $f_{\alpha}\stackrel{{\scriptstyle\tau_{\textrm{cc}}}}{{\longrightarrow}}f$ if and only if $f_{\alpha}\stackrel{{\scriptstyle d^{K}}}{{\longrightarrow}}f$ for every $K\in K(Y)$. (footnote [^8] Proof: ($\Rightarrow$): Assume $f_{\alpha}\stackrel{{\scriptstyle\tau_{\textrm{cc}}}}{{\longrightarrow}}f$. Fix any $K\in K(Y)$. Then for any collection $\mathcal{O}=\{O_{y}\ni f(y):y\in K\}$ of open sets in $X$, there is $\alpha^{\mathcal{O}}$ such that $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O_{y}$ (for all $y\in K$). In particular, for any $\varepsilon>0$, with $\mathcal{O}_{\varepsilon}:=\{O_{y,\varepsilon}:=B_{\varepsilon}^{K}(f(y))~|~y\in K\}$, $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\mathcal{O}_{\varepsilon}}}\subset O_{y,\varepsilon}~\forall y\in K$ implies $d^{K}(f_{\alpha},f)=\sup_{y\in K}d(f_{\alpha}(y),f(y))<\varepsilon~~~~\forall\alpha\geq\alpha^{\varepsilon}:=\alpha^{\mathcal{O}_{\varepsilon}}$. ($\Leftarrow$): Assume $f_{\alpha}\stackrel{{\scriptstyle d^{K}}}{{\longrightarrow}}f$ in $C(Y,X)$ for all $K\in K(Y)$. Fix $K\in K(Y)$. For any $\varepsilon>0$, there is $\alpha^{\varepsilon}$ such that $d^{K}(f_{\alpha},f)<\varepsilon$ for all $\alpha\geq\alpha^{\varepsilon}$. Therefore, $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\varepsilon}}\subset B_{\varepsilon}(f(y))$ for all $\in K$. Consider any collection $\mathcal{O}=\{O_{y}\ni f(y):y\in K\}$ of open sets in $X$. Let $B_{r_{y}}(f(y))\subset O_{y}$ for each $y\in K$. Then by the compactness of $f(K)$, we can choose $r>0$ (independent of $y$) such that $B_{r}(f(y))\subset O_{y}$ for each $y\in K$. Hence, with $\varepsilon<r$ and $\alpha^{\mathcal{O}}:=\alpha^{\varepsilon}$, we get $\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\mathcal{O}}}=\{f_{\alpha}(y)\}_{\alpha\geq\alpha^{\varepsilon}}\subset B_{\varepsilon}(f(y))\subset B_{r}(f(y))\subset O_{y}$ for all $y\in K$. ).
The topology of local uniform convergence $\tau_{luc}$ on $\mathcal{F}\subset X^{Y}$ is the topology with subbase
$\mathcal{S}\mathcal{B}_{luc}:=\left\{B^{d^{K}}_{r}(f):=\{g\in\mathcal{F}:d^{K}(f,g)<r\}~|~r>0,f\in\mathcal{F},K\in K(Y)\right\}.$
Note 6.
By Note 5 above: (i) On any $\mathcal{F}\subset X^{Y}$,   $\tau_{\textrm{cc}}\supset\tau_{luc}:=\bigcap\{\tau_{d^{K}}:K\in K(Y)\}$,   where a set $O\subset\mathcal{F}$ is $\tau_{luc}$-open (resp., $C\subset\mathcal{F}$ is $\tau_{luc}$-closed) $\iff$ $d^{K}$-open (resp., $d^{K}$-closed) for all $K\in K(Y)$. (ii) On continuous maps $\mathcal{F}:=C(Y,X)\subset X^{Y}$, $\tau_{\textrm{cc}}=\tau_{luc}$.
Definition 13 (An interpretation of the Vietoris topology).
Let $X$ be a space and $A_{\alpha},A\in\mathcal{P}^{\ast}(X)$. We say $A_{\alpha}$ centrally-converges (resp., marginally-converges) to $A$ in $X$, written $A_{\alpha}\stackrel{{\scriptstyle+}}{{\longrightarrow}}A$ (resp., $A_{\alpha}\stackrel{{\scriptstyle-}}{{\longrightarrow}}A$), if for any open set $O\subset X$ containing $A$ (resp., meeting $A$), there exists $\alpha_{+}^{O}$ (resp., $\alpha_{-}^{O}$) such that
$A_{\alpha}\subset O~\textrm{~}~\forall\alpha\geq\alpha_{+}^{O}~\textrm{~}~(\textrm{resp.,~}~A_{\alpha}\cap O\neq\emptyset~{~}~\forall\alpha\geq\alpha_{-}^{O}),$
that is, if $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{+}}}{{\longrightarrow}}A$ (resp., $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}A$). A set $\mathcal{C}\subset P^{\ast}(X)$ is centrally-closed (resp., marginally-closed) if for any net $A_{\alpha}\in\mathcal{C}$ such that $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{+}}}{{\longrightarrow}}A$ (resp., $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}A$), we have $A\in\mathcal{C}$. The upper Vietoris topology $\tau_{v}^{+}$ (resp., lower Vietoris topology $\tau_{v}^{-}$) of $\mathcal{P}^{\ast}(X)$ is the topology whose closed sets are the centrally-closed (resp., marginally-closed) subsets of $\mathcal{P}^{\ast}(X)$. Noting that $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}}}{{\longrightarrow}}A$ if and only if $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{+}}}{{\longrightarrow}}A$ and $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}A$, the Vietoris topology $\tau_{v}$ of $\mathcal{P}^{\ast}(X)$ is the topology whose closed sets are those subsets of $\mathcal{P}^{\ast}(X)$ that are each both centrally-closed and marginally-closed.
Lemma 4.1.
Let $X=(X,\mathcal{U})$ be a chained-space, $Y$ a space, $Z\subset Y$, and $f_{\alpha}\in X^{Y}$ a net. (i) If $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$, then $f_{\alpha}(Z)\stackrel{{\scriptstyle\tau_{v}^{+}}}{{\longrightarrow}}f(Z)$. (ii) If $f_{\alpha}|_{Z}\stackrel{{\scriptstyle\tau_{p}}}{{\longrightarrow}}f|_{Z}$, then $f_{\alpha}(Z)\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}f(Z)$ (iff $\overline{f_{\alpha}(Z)}\stackrel{{\scriptstyle\tau_{v}^{-}}}{{\longrightarrow}}\overline{f(Z)}$). (iii) If $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$, then $f_{\alpha}(Z)\stackrel{{\scriptstyle\tau_{v}}}{{\longrightarrow}}f(Z)$.
Proof.
(i) Pick any open neighborhood $O\in\mathcal{U}_{\lambda}$ of $f(Z)$. Consider the system of open sets $\mathcal{O}:=\{O_{y}=O\ni f(z)\}_{z\in Z}=\{O\}$. Then there exists $\alpha^{\mathcal{O}}$ such that for each $z\in Z$, $\{f_{\alpha}(z)\}_{\alpha\geq\alpha^{\mathcal{O}}}\subset O$, which implies $\bigcup_{\alpha\geq\alpha^{\mathcal{O}}}f_{\alpha}(Z)\subset O$. (ii) Let $O\in\mathcal{U}_{\lambda}$ be an open set such that $O\cap f(Z)\neq\emptyset$ (i.e., $f(Z)\in O^{-}$). Then some $f(z)\in O$. Since $f_{\alpha}(z)\rightarrow f(z)$, there is $\alpha_{z}^{O}$ such that $\{f_{\alpha}(z)\}_{\alpha\geq\alpha_{z}^{O}}\subset O$, i.e., $f_{\alpha}(Z)\cap O\neq\emptyset$ for all $\alpha\geq\alpha_{z}^{O}$. (iii) By Note 2 of Definition 12, $f_{\alpha}\stackrel{{\scriptstyle u|_{Z}}}{{\longrightarrow}}f$ implies $f_{\alpha}|_{Z}\stackrel{{\scriptstyle\tau_{p}}}{{\longrightarrow}}f|_{Z}$. So, the conclusion follows by (i) and (ii). ∎
Theorem 4.2 (Construction of an swrc-topology).
Let $X=(X,\mathcal{U})$ be a chained-space, $Y$ a space, and $\mathcal{Z}\subset\mathcal{P}^{\ast}(Y)$ a cover of $Y$ (i.e., $\bigcup\mathcal{Z}=Y$). Let $\mathcal{F}\subset X^{Y}$ be a $q$-full (and finitely $q$-stable) subspace such that (i) $\overline{f(Z)}=f(Z)$ for all $\{f\in\mathcal{F},Z\in\mathcal{Z}\}$, and (ii) $(\mathcal{F},\tau)$ is an rc-space with $\tau_{uc|_{\mathcal{Z}}}\subset\tau$ (for example, we could take $\tau:=\tau_{uc}$). Then the topology $\tau_{\pi}:=\tau_{uc|_{\mathcal{Z}}}$ on $\mathcal{F}\subset X^{Y}$ is a $\langle\tau_{p}\cup q^{-1}(\tau_{v})\rangle$-compatible wrc-topology and, moreover, $q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau_{v})$ is continuous (i.e., $\tau_{v}\subset\tau_{\pi q}$ in $Cl_{Y}(X)$ ).
NB: When $Y$ is finite, it suffices (by Theorem 3.3) to replace $\tau_{uc|_{\mathcal{Z}}}$ with $\tau_{p}$, in which case, Tychonoff’s product theorem ensures that (i) and (ii) are no longer needed.
Proof.
We need to verify the necessary requirement for $\tau_{\pi}$ in Definition 7. By Note 4 of Definition 12 and the equality $\bigcup\mathcal{Z}=Y$, we have $\tau_{p}\subset\tau_{uc|_{\mathcal{Z}}}$. Since $\mathcal{F}$ satisfies (i), we also have $q^{-1}(\tau_{v})\subset\tau_{uc|_{\mathcal{Z}}}$ by Lemma 4.1(iii). Therefore $\langle\tau_{p}\cup q^{-1}(\tau_{v})\rangle\subset\tau_{uc|_{\mathcal{Z}}}$.
  1. (1)
    $\tau_{\pi}$ is an rc-topology (hence a wrc-topology) on $\mathcal{F}$: Indeed, for any cr-net $\{f_{\alpha}\}_{\alpha\in\mathcal{A}}\subset\mathcal{F}$, the rc-space $(\mathcal{F},\tau)$ gives a subnet $f_{\alpha(\beta)}\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}f\in\mathcal{F}$. So $f_{\alpha(\beta)}\stackrel{{\scriptstyle\tau_{uc|_{\mathcal{Z}}}}}{{\longrightarrow}}f\in\mathcal{F}$ (since $\tau_{uc|_{\mathcal{Z}}}\subset\tau$).
    NB: It is clear that if $Y$ is finite, in which case $\tau_{uc|_{\mathcal{Z}}}\stackrel{{\scriptstyle(\ast)}}{{=}}\tau_{p}\stackrel{{\scriptstyle(\ast)}}{{=}}\langle\tau_{p}\cup q^{-1}(\tau_{v})\rangle=\langle\tau_{p}\cup q^{-1}(\tau_{v}^{+})\rangle$ (where the equalities ($\ast$) hold because pointwise convergence automatically implies uniform convergence, which in turn implies $q^{-1}(\tau_{v})$-convergence by Lemma 4.1(iii)), then the conclusion no longer requires (i) and (ii).
  2. (2)
    $q:(\mathcal{F},\tau_{\pi})\rightarrow(Cl_{Y}(X),\tau_{v})$ is continuous (hence continuous on $K\mathcal{F}(Y,X)$), since by construction $q^{-1}(\tau_{v})\subset\tau_{\pi}$ (hence $\tau_{v}\subset\tau_{\pi q}$).
Metrization of compact-subset hyperspaces
The existence of a swrc-topology on a $q$-full $\mathcal{F}\subset X^{Y}$ (say as in Theorem 4.2) allows us to concretely establish metrizability (in Theorems 4.6 and 4.9) of indexed compact-subset hyperspaces $K_{Y}(X)\subset K(X)$ of a metrizable space $X$.
Lemma 4.3 (Compact union I).
Let $X$ be a space and $Y$ a set. If $\mathcal{A}\subset(K_{Y}(X),\tau_{v})$ is compact, then so is $K:=\bigcup_{A\in\mathcal{A}}A\subset X$. Hence, if $\mathcal{B}\subset(K_{Y}(X),\tau_{v})$ is contained in a compact subset of $(K_{Y}(X),\tau_{v})$, then $L:=\bigcup_{B\in\mathcal{B}}B$ is contained in a compact subset of $X$.
Proof.
Consider a net $x_{\alpha}\in K$, and let $x_{\alpha}\in A_{\alpha}\in\mathcal{A}$. By the compactness of $\mathcal{A}$, let a subnet $A_{\alpha(\beta)}\stackrel{{\scriptstyle\tau_{v}}}{{\longrightarrow}}A_{0}\in\mathcal{A}$ (i.e., for any $\mathcal{O}:=[\{O_{1},...,O_{n}\}]_{v}\ni A_{0}$, some tail $\{A_{\alpha(\beta)\geq\alpha(\beta^{\mathcal{O}})}\}\subset[\{O_{1},...,O_{n}\}]_{v}$). We will show there is $x\in A_{0}$ such that $x_{\alpha(\beta)}\rightarrow x$. On the contrary, suppose that for all $x\in A_{0}$, $x_{\alpha(\beta)}\not\!\!\longrightarrow x$, i.e., there exists an open set $O_{x}\ni x$ and a subnet $\{x_{\alpha\circ\beta_{x}(\gamma)}\}$ that avoids $O_{x}$. In particular, since $n$ is finite, (a tail of) some subnet $\{x_{\alpha\circ\beta(\gamma)}\}$ avoids $O_{1},...,O_{n}$. But $\{A_{\alpha\circ\beta(\gamma)\geq\alpha\circ\beta(\gamma^{\mathcal{O}})}\}\subset[\{O_{1},...,O_{n}\}]_{v}$ implies $\{x_{\alpha\circ\beta(\gamma)\geq\alpha\circ\beta(\gamma^{\mathcal{O}})}\}\subset\bigcup_{i}O_{i}$ (a contradiction). (footnote [^9] Alternative proof: Let $\{O_{i}:i\in I\}$ be an open cover of $K$ (hence also an open cover of each $A\in\mathcal{A}$) in $X$. For each $A\in\mathcal{A}$, let $\{O^{A}_{i_{j}}:j\in I_{A}\}\subset\{O_{i}:i\in I\}$ be a finite subcover of AA (where wlog $A\cap O_{i_{j}}^{A}\neq\emptyset$ for all $j$). Then $A\in\langle O^{A}_{i_{j}}:j\in I_{A}\rangle_{v}$, and so $\big\{[\{O^{A}_{i_{j}}:j\in I_{A}\}]_{v}\big\}_{A\in\mathcal{A}}$ is a $\tau_{v}$-open cover of $\mathcal{A}$, which therefore has a finite $\tau_{v}$-subscover $\big\{[\{O^{A}_{i_{j}}:j\in I_{A}\}]_{v}\big\}_{A\in F}$, for some finite set $F\subset\mathcal{A}$. Hence $\{O^{A}_{i_{j}}:j\in I_{A},A\in F\}$ is a finite subcover of $K$ in $X$.). ∎
When $X$ is a metric space (in which case $(K(X),\tau_{v})\cong(K(X),\tau_{d_{H}})$ by Lemma 4.5), the above result becomes equivalent to the following.
Lemma 4.4 (Compact union II: [akofor2024, Lemma 3.5(ii)]).
Let $X=(X,d)$ be a metric space and $Y$ a set. If $\mathcal{C}\subset\big(K_{Y}(X),d_{H}\big)$ is compact, then $K:=\bigcup_{C\in\mathcal{C}}C\subset X$ is compact.
We review the proof of the following well known result.
Lemma 4.5 ([IllanNadl1999, Theorem 3.1]).
If $X$ is a metrizable space, then so is $(K(X),\tau_{v})$. Moreover, if $X\cong(X,d)$, then $(K(X),\tau_{v})\cong(K(X),d_{H})$. Conversely, if $X$ is $T_{1}$ (i.e., all singletons of $X$ are closed) and $K(X)$ is metrizable, then $X$ is metrizable (as a subspace of a metrizable space).
Proof.
Let $X=(X,d)$. (i) Showing $\tau_{v}\subset\tau_{d_{H}}$ in $K(X)$: Let $O\subset X$ be open. We need to show $O^{+}$ and $O^{-}$ are in $\tau_{d_{H}}$. First, let $A_{n}\in(O^{+})^{c}=\{K\in K(X):K\cap O^{c}\neq\emptyset\}=(O^{c})^{-}$ such that $A_{n}\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}A$. Then with $a_{n}\in A_{n}\cap O^{c}\subset A\cup\bigcup_{n}A_{n}$ (a compact set), let $d(a_{f(n)},e)\rightarrow 0$, where $e\in O^{c}$ (a closed set). Then
$d(e,A)$ $=\inf_{a\in A}d(e,a)\leq\inf_{a\in A}[d(e,a_{f(n)})+d(a_{f(n)},a)]=d(e,a_{f(n)})+d(a_{f(n)},A)$
$\leq d(e,a_{f(n)})+d_{H}(A_{f(n)},A)\rightarrow 0,$
$~~\Rightarrow~~e\in A~~\Rightarrow~~O^{c}\cap A\neq\emptyset,~~\Rightarrow~~A\in(O^{+})^{c},$
and so $(O^{+})^{c}$ is $d_{H}$-closed, i.e., $O^{+}$ is $d_{H}$-open.
Second, let $A_{n}\in(O^{-})^{c}=\{K\in K(X):K\subset O^{c}\}=(O^{c})^{+}$ such that $A_{n}\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}A$. Fix $a\in A$. Then, for some $a_{n}=a_{n}(a)\in A_{n}$, we have $d(a,a_{n})=d(a,A_{n})\leq d_{H}(A,A_{n})\rightarrow 0$. Since $a_{n}\in A\cup\bigcup_{n}A_{n}$ (a compact set), let $d(a_{n},e)\rightarrow 0$ for some $e\in O^{c}$ (since $A_{n}\subset O^{c}$ and $O^{c}$ is a closed set). Then
$d(a,e)\leq d(a,a_{n})+d(a_{n},e)\rightarrow 0,~~\Rightarrow~~a=e,~~\Rightarrow~~A\subset O^{c},$
and so $(O^{-})^{c}$ is $d_{H}$-closed (i.e., $O^{-}$ is $d_{H}$-open).
(ii) Showing $\tau_{d_{H}}\subset\tau_{v}$ in $K(X)$: Let $\mathcal{C}\subset K(X)$ be a $d_{H}$-closed set. Let $A_{\alpha}\in\mathcal{C}$ such that $A_{\alpha}\stackrel{{\scriptstyle\tau_{v}}}{{\longrightarrow}}A$. We need to show $A\in\mathcal{C}$. Suppose $A\in\mathcal{C}^{c}$. For any $\mathcal{O}:=[O_{1},...,O_{n}]_{v}\ni A$, some tail $\{A_{\alpha\geq\alpha^{\mathcal{O}}}\}\subset[O_{1},...,O_{n}]_{v}$. With $\varepsilon^{\mathcal{O}}:=\max_{i}\textrm{diam}(O_{i})$, we get  $\{A_{\alpha\geq\alpha^{\mathcal{O}}}\}\subset\overline{B_{\varepsilon^{\mathcal{O}}}^{d_{H}}(A)}$,  since for $\alpha\geq\alpha^{\mathcal{O}}$,
$\textstyle d_{H}(A,A_{\alpha})$ $=\max_{A\leftrightarrow A_{\alpha}}\max_{a\in A}d(a,A_{\alpha})\leq\max_{A\leftrightarrow A_{\alpha}}\max_{a\in A}d(a,A_{\alpha}\cap O_{i_{a}})|_{a\in O_{i_{a}}}\leq\varepsilon^{\mathcal{O}}.$
Since $A\in\mathcal{C}^{c}$ (a $d_{H}$-open set), some $B_{r}^{d_{H}}(A)\subset\mathcal{C}^{c}$. So, by choosing $\mathcal{O}$ such that $\varepsilon^{\mathcal{O}}<r$ (which is possible by the compactness of $A$), we get $\{A_{\alpha\geq\alpha^{\mathcal{O}}}\}\subset\mathcal{C}^{c}$ (a contradiction). ∎
Theorem 4.6.
If $X=(X,d)$ is a metric space and $Y$ a finite set, then  $(Cl_{Y}(X),\tau_{pq})=(Cl_{Y}(X),\tau_{v})=(Cl_{Y}(X),d_{H})$.
Proof.
Since $Cl_{Y}(X)=FS_{|Y|}(X)\subset K(X)$, it follows from Lemma 4.5 that $(Cl_{Y}(X),\tau_{v})=(Cl_{Y}(X),d_{H})$”, where by Theorem 3.3, $\tau_{pq}=\tau_{v}$ in $Cl_{Y}(X)$. ∎
Theorem 4.6 is metrization of the quotient of the swrc-topology $(X^{Y},\tau_{p})$ due to a finite $Y$. Metrization of the quotient of a general $q$-full swrc-space $(\mathcal{F},\tau_{\pi})\subset X^{Y}$ is given by Theorem 4.9.
Lemma 4.7.
Let $X$ be a set, $(Y,\tau)$ a space, $f:X\rightarrow Y$ a map, and $y_{\alpha}\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}y$ a convergent net in $Y$. If $x_{\alpha}\in f^{-1}(y_{\alpha})$ and $x\in f^{-1}(y)$, then  $x_{\alpha}\stackrel{{\scriptstyle f^{-1}(\tau)}}{{\longrightarrow}}x$ in $X$ (where $f^{-1}(\tau):=\{f^{-1}(U):U\in\tau\}$). (footnote [^10] Since $f:(X,f^{-1}(\tau))\rightarrow(Y,\tau)$ is continuous, for every convergent net $x_{\alpha}\stackrel{{\scriptstyle f^{-1}(\tau)}}{{\longrightarrow}}x$ in $X$, we have $f(x_{\alpha})\stackrel{{\scriptstyle\tau}}{{\longrightarrow}}f(x)$ in $Y$. ).
Proof.
Fix any $x_{\alpha}\in f^{-1}(y_{\alpha})$ and any $x\in f^{-1}(y)$. Let $O\ni x$ be an $f^{-1}(\tau)$-open neighborhood of $x$ in $X$, i.e., $x\in O=f^{-1}(U)$ for an open set $U\subset Y$. Then $y=f(x)\in U$ and so some tail $\{y_{\alpha}\}_{\alpha\geq\alpha^{U}}\subset U$. Therefore, $\{x_{\alpha}\}_{\alpha\geq\alpha^{U}}\subset f^{-1}(\{y_{\alpha}\}_{\alpha\geq\alpha^{U}})\subset f^{-1}(U)=O$. ∎
Theorem 4.8.
Let $X=(X,d)$ be a metric space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$ a $q$-full subset. Then the space $\big(K\mathcal{F}(Y,X),q^{-1}(\tau_{v})\big)$ of precompact image maps $K\mathcal{F}(Y,X):=\mathcal{F}\cap q^{-1}(K_{Y}(X))$ in $\mathcal{F}$ is pseudometrized by
$d^{q}_{H}(f,g):=d_{H}\left(q(f),q(g)\right).$
That is, $(K\mathcal{F}(Y,X),q^{-1}(\tau_{v}))\cong\big(K\mathcal{F}(Y,X),q^{-1}(\tau_{d_{H}})\big)$.
Proof.
This is precisely the proof of Lemma 4.5, with $K(X)$ replaced by $K_{Y}(X)$, along with the following basic observations (for an open set $O\subset X$): (1) $A=q(f)\in O^{\pm}$ $\iff$ $f\in q^{-1}(O^{\pm})$ in $\mathcal{F}$. (2) $B=q(g)\in B_{r}^{d_{H}}\left(q(f)\right)$ $\iff$ $g\in B_{r}^{d^{q}_{H}}(f)$ in $\mathcal{F}$. ∎
Remark.
In Theorem 4.8, if the continuous maps $C(Y,X)$ in particular satisfy $q(C(Y,X))=q(X^{Y})$, then with  $\mathcal{F}:=C(Y,X)$, the precompact image maps  $K\mathcal{F}(Y,X):=\mathcal{F}\cap q^{-1}(K_{Y}(X))$  can be replaced with the precompact image continuous maps  $KC(Y,X):=C(Y,X)\cap q^{-1}(K_{Y}(X))$.
Theorem 4.9 (Metrization of compact-subset hyperspaces).
Let $X=(X,d)$ be a metric space, $Y$ a set, and $(\mathcal{F},\tau_{\pi})\subset X^{Y}$ a $q$-full swrc-space. Then $(K_{Y}(X),\tau_{\pi q})\cong(K_{Y}(X),d_{H})\cong(K_{Y}(X),\tau_{v})$.
Proof.
By the continuity of the map
[…]
the $\tau_{\pi q}$-topology of $K_{Y}(X)$ contains the $d_{H}$-topology (i.e., $\tau_{d_{H}}\subset\tau_{\pi q}$). This shows, among other things, that $(K_{Y}(X),\tau_{\pi q})$ is Hausdorff, since $\tau_{d_{H}}$ is Hausdorff. In $K_{Y}(X)$, we also have $\tau_{\pi q}\subset\tau_{d_{H}}$, as shown next.
Let $\mathcal{C}\subset K_{Y}(X)$ be a $\tau_{\pi q}$-closed set. Consider a sequence $\left\{q(f_{n})\right\}_{n}\subset\mathcal{C}$ such that $q(f_{n})\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}q(f)$ for some $q(f)\in BCl_{Y}(X)$. We want to show that $q(f)\in\mathcal{C}$ (in which case $\mathcal{C}$ is also $d_{H}$-closed). In $\mathcal{F}$, let $g_{n}\in q^{-1}\left(q(f_{n})\right)\subset q^{-1}(\mathcal{C})$, i.e., $q(g_{n})=q(f_{n})$ (footnote [^11] A natural choice here is $g_{n}:=f_{n}$. So, when $Y$ is finite, it is enough to take $\tau_{\pi}:=\tau_{p}$ and thereby automatically get a pointwise convergent subnet of the crcr-sequence $\{g_{n}\}$ by Tychonoff’s product theorem (thereby obtaining an alternative proof of Theorem 4.6 without Theorem 3.3).).
Since $\left\{q(f)\right\}\cup\left\{q(f_{n})\right\}_{n}\subset\big(K_{Y}(X),d_{H}\big)$ is compact, $K=q(f)\cup\bigcup_{n}q(f_{n})\subset(X,d)$ is compact by Lemma 4.4. So, by the definition of $\tau_{\pi}$, the cr-sequence $\{g_{n}:Y\rightarrow q(f_{n})\subset K\subset(X,d)\}_{n}$ has a $\tau_{\pi}$-convergent subnet $\{g_{n(\alpha)}\}\stackrel{{\scriptstyle\tau_{\pi}}}{{\longrightarrow}}h$, where $h\in q^{-1}(\mathcal{C})$ since $q^{-1}(\mathcal{C})$ is $\tau_{\pi}$-closed. By the continuity of (4),
$q(g_{n(\alpha)})=q(f_{n(\alpha)})\stackrel{{\scriptstyle d_{H}}}{{\longrightarrow}}q(f)=q(h)\in\mathcal{C}.$
5. Quotient-realization of hyperspace topologies
When $Y$ is not finite, it is not clear whether or not $\mathcal{F}\subset X^{Y}$ always admits a swrc-topology, even though we have seen in Theorem 4.2 that a swrc-topology can exist under special conditions (which allow us to obtain a swrc-topology from a $\tau_{p}$-compatible topology with a $\tau_{v}$-compatible quotient). We will prove the existence of a $\tau_{p}$-compatible topology with $\tau_{v}$ as its quotient in Theorem 5.1.
In this section, unless said otherwise, let $X$ be a space, $Y$ a set, and $\mathcal{F}\subset X^{Y}$ a $q$-full subset.
Definition 14 ($q$-lifts of subset hyperspace topologies).
Let $\tau$ be a topology on $Cl(X)$. A $q$-lift of $\tau$ is any topology $\widetilde{\tau}$ on $\mathcal{F}$ such that $\widetilde{\tau}_{q}=\tau$ in $Cl_{Y}(X)$. We show in Theorem 5.1 that if $\tau$ is $\tau_{v}$-compatible, then a $\tau_{p}$-compatible $q$-lift $\widetilde{\tau}$ of $\tau$ can be chosen and the choice is maximal with respect to subsets of a $q$-full finitely $q$-stable $\mathcal{F}\subset X^{Y}$.
Remark (Existence of $q$-lifts, Smallest $q$-lift).
Observe that given any topology $\tau_{0}$ on $Cl(X)$, the topology $q^{-1}(\tau_{0}):=\{q^{-1}(B):B\in\tau_{0}\}$ on $\mathcal{F}$ is the smallest $q$-lift of $\tau_{0}$ (where we note that $q:\big(\mathcal{F},q^{-1}(\tau_{0})\big)\rightarrow(Cl_{Y}(X),\tau_{0})$ is an open continuous map, hence an open quotient map). That is, $q^{-1}(\tau_{0})\subset\widetilde{\tau}_{0}$ for every $q$-lift $\widetilde{\tau}_{0}$ of $\tau_{0}$. So,  $q^{-1}(\tau_{0})=\bigcap\{\textrm{$q$-lifts $\widetilde{\tau}_{0}$ of $\tau_{0}$}\}$.
Remark (Temporary notation for the proof of Theorem 5.1).
Let $X$ be a space, $Y$ a set, $\mathcal{I},\mathcal{J}\subset X^{Y}$ subsets, $\tau_{i}$ any topology on $\mathcal{I}$, and $\tau_{j}$ any topology on $\mathcal{J}$. Consider the map $q:X^{Y}\rightarrow Cl_{Y}(X)$.
  1. (1)
    As usual, in any space $(X^{Y},\tau)$, with $\tau\cap\mathcal{I}:=\{A\cap\mathcal{I}:A\in\tau\}$, we write the subspace $(\mathcal{I},\tau\cap\mathcal{I})\subset(X^{Y},\tau)$ simply as $(\mathcal{I},\tau)\subset(X^{Y},\tau)$ or as $\mathcal{I}\subset(X^{Y},\tau)$.
  2. (2)
    In the restriction $q:(\mathcal{I},\tau_{i})\subset X^{Y}\rightarrow(q(\mathcal{I}),\tau_{iq})\subset Cl_{Y}(X)$, if $B\subset Cl_{Y}(X)$, then we write $B\in\tau_{iq}$” to mean $B\cap q(\mathcal{I})\in\tau_{iq}$”.
  3. (3)
    In the restrictions $q:(\mathcal{I},\tau_{i})\rightarrow(q(\mathcal{I}),\tau_{iq})$ and $q:(\mathcal{J},\tau_{j})\rightarrow(q(\mathcal{J}),\tau_{jq})$, we write $\tau_{iq}\subset\tau_{jq}$” to mean $\forall B\subset Cl_{Y}(X)$, $B\in\tau_{iq}$ $\Rightarrow$ $B\in\tau_{jq}$ (i.e., $B\cap q(\mathcal{I})\in\tau_{iq}$ $\Rightarrow$ $B\cap q(\mathcal{J})\in\tau_{jq}$)”.
Theorem 5.1 (Existence of $\tau_{p}$-compatible $q$-lifts).
Let $X$ be a space, $Y$ a set, $\mathcal{F}\subset X^{Y}$ a $q$-full finitely $q$-stable subset, and $\tau_{0}\supset\tau_{v}$ a topology on $Cl(X)$. Further suppose that some finite-subset hyperspace $FS_{n}(X)\subset(Cl_{Y}(X),\tau_{v})$ is closed. Then there exists a topology $\widetilde{\tau}_{0}\supset\tau_{p}$ on $\mathcal{F}$, which is maximal with respect to subsets of $\mathcal{F}$, such that $\tau_{0}\supset\widetilde{\tau}_{0q}\supset\tau_{v}$ in $Cl_{Y}(X)$.
Proof.
[…]
By Theorem 3.3(1) and Theorem 3.3’s Remark (1), $\mathcal{P}\neq\emptyset$ (because with $1\leq n\leq|Y|$, $\mathcal{A}:=\{f\in\mathcal{F}:|q(f)|\leq n\}=q^{-1}(FS_{n}(X))\cap\mathcal{F}$, and $\tau:=\tau_{p}$, we get $\tau_{q}=\tau_{v}$ in $q:(\mathcal{A},\tau)\rightarrow(q(\mathcal{A}),\tau_{q})\subset FS_{n}(X)$). Given a chain $\{(\mathcal{A}_{\lambda},\tau_{\lambda})\}_{\lambda\in\Lambda}$ in $\mathcal{P}$, let $\mathcal{A}:=\bigcup_{\lambda}\mathcal{A}_{\lambda}$ and $\tau:=\bigcup_{\lambda}\tau_{\lambda}$ (i.e., for any $A\subset\mathcal{A}$, we have $\in\tau \iff A𝒜$$\cap\mathcal{A}_{\lambda}\in\tau_{\lambda}$ for each $\lambda$). Then (i) $\mathcal{A}\subset\mathcal{F}$, (ii) $\tau_{p}\subset\tau$, and (iii) $\tau_{v}\subset\tau_{q}\subset\tau_{0}$ (since $\tau_{v}\subset\tau_{\lambda q}\subset\tau_{0}$ for each $\lambda$ and $\tau_{q}=\bigcap_{\lambda}\tau_{\lambda q}$). Here, $\tau_{q}=\bigcap_{\lambda}\tau_{\lambda q}$ is due to the following: Given $B\in Cl_{Y}(X)$, for each $\lambda$, $B\in\tau_{\lambda q}$ $\stackrel{{\scriptstyle\textrm{def.}}}{{\iff}}$ $q^{-1}(B)\cap\mathcal{A}_{\lambda}\in\tau_{\lambda}$, and so $B\in\tau_{q}$ $\iff$ $q^{-1}(B)\in\tau$ $\stackrel{{\scriptstyle\textrm{def.}}}{{\iff}}$ $q^{-1}(B)\cap\mathcal{A}_{\lambda}\in\tau_{\lambda}$ $\forall\lambda$ $\iff$ $B\in\tau_{\lambda q}$ $\forall\lambda$.
Therefore $(\mathcal{A},\tau)$ is an upper bound of the given chain in $\mathcal{P}$. By Zorn’s lemma, $\mathcal{P}$ has a maximal element $(\mathcal{A}^{\prime},\tau^{\prime})$. Suppose $\mathcal{A}^{\prime}\neq\mathcal{F}$. Let $f\in\mathcal{F}\backslash\mathcal{A}^{\prime}$, $\mathcal{A}^{\prime\prime}:=\mathcal{A}^{\prime}\cup\{f\}$, and $\tau^{\prime\prime}$ be the topology on $\mathcal{A}^{\prime\prime}$ given by $\tau^{\prime\prime}:=\{A\subset\mathcal{A}^{\prime\prime}:A\cap\mathcal{A}^{\prime}\in\tau^{\prime}\}=\tau^{\prime}\cup\big\{A\cup\{f\}:A\in\tau^{\prime}\big\}$ (footnote [^13] This is the superspace topology defined in the footnote(s) on page Remark.).
We claim that $(\mathcal{A}^{\prime},\tau^{\prime})<\big(\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big)\in\mathcal{P}$. Proof of claim: It is clear that $\mathcal{A}^{\prime}\subsetneq\mathcal{A}^{\prime\prime}$ and $\tau^{\prime}\subset\tau^{\prime\prime}$ (which implies $\tau^{\prime}_{q}\supset\tau_{q}^{\prime\prime}$ as before), and so $(\mathcal{A}^{\prime},\tau^{\prime})<\big(\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big)$. Next, $\big(\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big)\in\mathcal{P}$ follows from (i) $\mathcal{A}^{\prime\prime}\subset\mathcal{F}$, (ii) $\tau_{p}\subset\tau^{\prime\prime}$, and (iii) $\tau_{v}\subset\tau_{q}^{\prime}=\tau_{q}^{\prime\prime}\subset\tau_{0}$, since we also have $\tau_{q}^{\prime}\subset\tau_{q}^{\prime\prime}$ (because given $B\in Cl_{Y}(X)$, $B\in\tau_{q}^{\prime}$ $\iff$ $q^{-1}(B)\cap\mathcal{A}^{\prime}=(q^{-1}(B)\cap\mathcal{A}^{\prime\prime})\cap\mathcal{A}^{\prime}\in\tau^{\prime}$ $\stackrel{{\scriptstyle\textrm{construction of $\tau^{\prime\prime}$}}}{{\Longrightarrow}}$ $q^{-1}(B)\cap\mathcal{A}^{\prime\prime}\in\tau^{\prime\prime}$ $\iff$ $B\in\tau_{q}^{\prime\prime}$). This completes the proof of the claim.
But $(\mathcal{A}^{\prime},\tau^{\prime})<\big(\mathcal{A}^{\prime\prime},\tau^{\prime\prime}\big)\in\mathcal{P}$ contradicts maximality of $(\mathcal{A}^{\prime},\tau^{\prime})$ in $\mathcal{P}$. So, we can set $\widetilde{\tau}_{0}:=\tau^{\prime}$. ∎
Question 2.
In Theorem 5.1, when can we choose the topology $\widetilde{\tau}_{0}$ to be a rc-, wrc-, or swrc-topology?
6. Conclusion and questions
For a space $X$, we have seen (Theorem 3.3) that certain hyperspaces $\mathcal{J}\subset Cl(X)$ of $X$ can be described as quotients of function spaces $\mathcal{F}\subset X^{Y}$ in a natural way. Following this, we have discussed the concrete realization of certain preferred function space topologies (Theorem 4.2), the metrization of compact-subset hyperspaces (Theorems 4.6 and 4.9), and the existence of $\tau_{p}$-compatible function space topologies with $\tau_{v}$-compatible quotients (Theorem 5.1).
In addition to Questions 1 and 2, we have the following interesting questions.
Question 3.
Let $X,Y$ be spaces, $\mathcal{F}\subset X^{Y}$ a $q$-full subset, $Z\subset Y$, and $Cl_{Z}(X)\subset(Cl_{Y}(X),\tau)$ a $\tau$-closed subset, in which case a quotient map $q:(\mathcal{F},\widetilde{\tau})\rightarrow(Cl_{Y}(X),\tau)$ restricts to a quotient map $q:(\mathcal{F}_{Z},\widetilde{\tau})\subset X^{Y}\rightarrow(Cl_{Z}(X),\tau)$, where $\mathcal{F}_{Z}:=q^{-1}(Cl_{Z}(X))\cap\mathcal{F}$. If $X$ is a metric space, can we find a Lipschitz retraction
$(Cl_{Y}(X),\tau)\rightarrow(Cl_{Z}(X),\tau)?$
This question is of interest especially in the case where $Y$ is finite (see [akofor2019, akofor2020, AkoKov2021, BacKov2016, COR2021, Kovalev2016] and references therein).
Question 4.
Let $X$ be a space and $Y$ a set. From Definition 3, for which $Y$ (of smallest possible cardinality) does the equality $Cl_{Y}(X)=Cl(X)$ hold? Also, what is the cardinally-smallest $q$-full subset $\mathcal{F}\subset X^{Y}$?
Question 5 (Path representation problem).
Let $X$ be a space, $Y$ a set, $\mathcal{F}\subset X^{Y}$ a $q$-full subset, and $\tau_{\pi}$ a swrc-topology on $\mathcal{F}$. It is clear that a path $\eta:[0,1]\rightarrow(\mathcal{F},\tau_{\pi})$ gives a path $q\circ\eta:[0,1]\stackrel{{\scriptstyle\eta}}{{\longrightarrow}}\mathcal{F}\stackrel{{\scriptstyle q}}{{\longrightarrow}}(Cl_{Y}(X),\tau_{\pi q})$, since the composition of continuous maps is continuous. Conversely, (when) does every path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ come from (or lift as $\gamma=q\circ\eta$ to) a path $\eta:[0,1]\rightarrow(\mathcal{F},\tau_{\pi})$?
Whenever the answer to this question is positive, every path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ is expressible in the form
$\gamma(t)$ $=q\circ\eta(t)=q\big(\eta(t)(Y)\big)=cl_{X}\eta(t)(Y)=cl_{X}\{\eta(t)(y):y\in Y\}$
$\equiv cl_{X}{\{\gamma_{r}(t):r\in\Gamma\}},~~\textrm{~}\forall t\in[0,1],$
for a set of paths $\{\gamma_{r}:[0,1]\rightarrow X\}_{r\in\Gamma}$ in $X$. Of course, by the axiom of choice, any path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ can be written as
$\gamma=q\circ\eta:[0,1]\stackrel{{\scriptstyle\eta}}{{\longrightarrow}}\mathcal{F}\stackrel{{\scriptstyle q}}{{\longrightarrow}}(Cl_{Y}(X),\tau_{\pi q}),~t\mapsto q\circ\eta(t)=cl_{X}\big(\eta(t)(Y)\big)$
for a (not necessarily continuous) selection
$\eta:[0,1]\rightarrow(\mathcal{F},\tau_{\pi}),~\textrm{~}~t\mapsto\eta(t)~\in~q^{-1}(\gamma(t)).$
This implies every path $\gamma:[0,1]\rightarrow(Cl_{Y}(X),\tau_{\pi q})$ is naturally expressible in the form
$\gamma(t)=cl_{X}{\{\gamma_{r}(t):r\in\Gamma\}}\equiv cl_{X}\{\gamma_{y}(t):=\eta(t)(y)~|~y\in Y\},~\textrm{~}~\forall t\in[0,1],$
for (not necessarily continuous) maps $\{\gamma_{r}:[0,1]\rightarrow X\}_{r\in\Gamma}$ (which therefore need not be paths in $X$).
By the above paragraph, Question 5 is relevant to the “path representation” problem considered in [akofor2024] and therefore relevant to [akofor2024, Question 5.1], even though the current topology $\tau_{\pi q}$ on $BCl_{Y}(X)\subset Cl_{Y}(X)$ is in general not the same as the $d_{H}$-topology. Since $\gamma^{-1}(\mathcal{O})=\eta^{-1}(q^{-1}(\mathcal{O}))$ for any open set $\mathcal{O}\subset(Cl_{Y}(X),\tau_{\pi q})$, if $\tau_{\pi}=q^{-1}(\tau_{\pi q})$, then $\gamma$ is continuous iff $\eta$ is continuous (but in general, if we have proper containment $q^{-1}(\tau_{\pi q})\subsetneq\tau_{\pi}$, then $\gamma=q\circ\eta$ can be continuous even when $\eta$ is not continuous). Therefore, for $\gamma=q\circ\eta$ to be continuous, we only need a $q^{-1}(\tau_{\pi q})$-continuous (not necessarily a $\tau_{\pi}$-continuous) selection
$\eta:[0,1]\rightarrow(\mathcal{F},q^{-1}(\tau_{\pi q})),~\textrm{~}~t\mapsto\eta(t)~\in~q^{-1}(\gamma(t)).$
In particular, if $Y$ is finite and $X^{Y}$ is Hausdorff, then by [Vesely1991, Theorem 3.3], the continuous selection $\eta$ always exists, in which case, every path $\gamma:[0,1]\rightarrow Cl_{Y}(X)=FS_{|Y|}(X)$ has the form $\gamma(t)=\{\gamma_{r}(t):r\in\Gamma\}$, for paths $\gamma_{r}:[0,1]\rightarrow X$.
Acknowledgments
This manuscript has been improved using feedback from referees and editors of The Houston Journal of Mathematics.